### v1

# Negative Yields Positive: Unified Dual-Path Adapter for Vision-Language Models

Abstract
Recently, large-scale pre-trained Vision-Language Models (VLMs) have demonstrated great potential in learning open-world visual representations, and exhibit remarkable performance across a wide range of downstream tasks through efficient fine-tuning. In this work, we innovatively introduce the concept of dual learning into fine-tuning VLMs, i.e., we not only learn what an image is, but also what an image isn’t. Building on this concept, we introduce a novel DualAdapter approach to enable dual-path adaptation of VLMs from both positive and negative perspectives with only limited annotated samples. In the inference stage, our DualAdapter performs unified predictions by simultaneously conducting complementary positive selection and negative exclusion across target classes, thereby enhancing the overall recognition accuracy of VLMs in downstream tasks. Our extensive experimental results across 15 datasets validate that the proposed DualAdapter outperforms existing state-of-the-art methods on both few-shot learning and domain generalization tasks while achieving competitive computational efficiency. Code is available at https://github.com/zhangce01/DualAdapter.
Keywords:
Vision-Language Models CLIP Few-Shot Classification Domain Generalization Transfer Learning
1 Introduction
Large-scale pre-trained Vision-Language Models (VLMs), such as CLIP [48] and ALIGN [24], provide a new paradigm for multi-modal learning and open-world visual recognition [78, 8]. Recently, researchers have demonstrated the robustness and effectiveness of these VLMs in interpreting complex visual and textual inputs, as evidenced by their superior performance in a variety of vision-language tasks, e.g., image captioning [73, 21, 60], visual question answering [52, 25, 31], and other domains [7, 85, 54, 58].
Current VLMs excel in zero-shot inference [86, 15], but struggle in transferring to specific downstream tasks, primarily due to domain shifts and mismatches between pre-training and task-specific data [10, 42]. Additionally, the extensive number of parameters and demanding computational costs make it intractable for individual practitioners to re-train or fine-tune the entire models to enhance their task-specific performance [84, 12]. To address this, recent studies have focused on designing efficient fine-tuning strategies that maintain the core pre-trained encoders frozen while incorporating lightweight adapter [12, 79] modules or optimizing prompts[84, 83]. These methods enable effective task-specific knowledge extraction from a limited amount of annotated training samples, on top of the inherited foundational knowledge from pre-training.
Figure 1: Dual-path inference examples. We show four sample images from diverse recognition tasks, along with their similarities (scaled by 100) to positive/negative text descriptions after few-shot dual-path adaptation. Intuitively, an image is more likely to be classified into a specific class if it exhibits higher positive similarity and lower negative similarity. We can observe that while the positive branch alone may struggle to distinguish among some closely related classes, incorporating the negative path, which eliminates certain incorrect classes, enhances its ability to accurately identify the true class.
In this work, we introduce the innovative concept of dual inference in the context of fine-tuning VLMs. Take, for instance, a basic binary pet classification task: distinguishing cats from dogs. To identify an image as a cat, we can approach the task from both the positive and negative perspectives: one might directly identify the image as a cat (selection), or alternatively, one can also first determine that it is not a dog (exclusion), thereby indirectly concluding it’s a cat. These approaches, namely, positive selection and negative exclusion, complement each other and collectively enhance the overall classification accuracy in general classification tasks.
Building on this concept, we introduce a novel DualAdapter approach for few-shot adaptation of VLMs that conducts positive selection and negative exclusion across a set of target classes. While CLIP typically performs classification tasks using prompts such as “A photo of a {CLASS}”, we recognize its ability to understand negation, employing prompts like “A photo of no {CLASS}”. Utilizing this capability, we design our DualAdapter to adapt VLMs from a dual perspective: we propose to learn not only what an image is but also what an image isn’t by matching the input image feature with the positive and negative text features. Intuitively, our approach first determines a set of likely candidate classes through image classification but, in a second step, refines these classification logits by asking the inverse question of what an image isn’t across the entire set of target classes. In Fig. 1, we showcase several dual-path inference examples to illustrate that when positive selection struggles with distinguishing between similar classes, employing negative exclusion can provide further insights to improve recognition accuracy.
Moreover, in DualAdapter, we symmetrically apply dual-path adaptation to the visual modality, treating image features of a specific class as positive and those from all other classes as negative. Recognizing the variability in image quality and representativeness, where some image features are noisy and not equally representative of their respective classes, we implement an unsupervised similarity-based label refinement technique. This method assigns confidence scores to each image, reducing the impact of outliers or less representative samples during few-shot adaptation, ensuring more accurate and reliable classification.
To empirically validate DualAdapter, we perform thorough evaluations on few-shot learning and domain generalization tasks using 15 diverse recognition datasets. These experimental results demonstrate the effectiveness of DualAdapter in adapting VLMs for downstream tasks and verify its superior robustness to distribution shifts. Additionally, our efficiency benchmarks indicate that DualAdapter achieves competitive efficiency compared to other state-of-the-art methods.
In summary, our primary contributions are listed as follows:
  • $\bullet$
    We explore and exploit the negative inference capabilities of VLMs, and introduce a unique dual-path adaptation approach (i.e., DualAdapter) for CLIP.
  • $\bullet$
    To mitigate the impact of noisy samples in few-shot adaptation, we introduce a novel unsupervised similarity-based label refinement technique.
  • $\bullet$
    Through extensive experiments, we demonstrate that our DualAdapter outperforms other state-of-the-art methods in both adaptation performance and generalizability across 15 diverse recognition datasets.
2 Related Work
Vision-Language Models. In recent years, extensive Vision-Language Models (VLMs) have been developed to bridge the vision and language modalities through large-scale pre-training [48, 24]. Generally, these pre-trained VLMs can be classified into four categories based on their pre-training objectives: masked language modeling [37, 30, 27], masked region prediction [37, 30, 59], image-text matching [30, 27, 61, 31], and contrastive learning [48, 24, 74, 77, 33]. Researchers have demonstrated that these pre-trained VLMs exhibit remarkable open-world visual understanding capabilities [8], achieving state-of-the-art performance across a range of open-world tasks, such as zero-shot learning [48, 51, 41], open-world segmentation [34, 72, 49], and open-world detection [53, 16, 9]. In this work, we mainly focus on contrastive-based VLMs, especially CLIP [48], which has become the current mainstream approach in this field. Nonetheless, we believe that the core concept of this work is adaptable and can be extended to other VLMs as well.
Efficient Fine-Tuning for VLMs. Given the substantial size of the VLMs, recent research efforts [84, 12, 79, 80] are focusing on the development of lightweight fine-tuning techniques to efficiently adapt VLMs for downstream visual tasks. These methods can generally be divided into two primary categories: prompt-based learning and adapter-style fine-tuning. Specifically, prompt-based learning methods [84, 83, 2, 62] aim to optimize the input prompts from downstream data, while adapter-style fine-tuning approaches [12, 79, 64, 75, 88] directly tune the extracted visual and textual representations. Our proposed DualAdapter follows the line of adapter-style fine-tuning, but it distinguishes itself from previous work [79, 75] by incorporating negative adapters for the first time. Moreover, we empirically demonstrate that leveraging negative cues from CLIP can effectively improve both few-shot classification performance and generalization capability.
Negative Learning. Given that obtaining typical positive labels (which indicate the categories an image belongs to) can be costly and labor-intensive, researchers have proposed an alternative approach known as the indirect negative learning [23, 76] paradigm. This method focuses on learning from easily accessible negative/complementary labels, which specify the categories to which an image does not belong. In recent years, negative learning has been effectively applied to various vision applications, such as image recognition [23, 28, 13], few-shot learning [22, 69], and semantic segmentation [67, 47]. In this work, empowered by the strong image-text association capabilities of the advanced pre-trained CLIP [48] model, we enable negative learning within CLIP without the need for explicit negative labels. Instead, we design negative prompts and negative image samples to guide the CLIP model to predict from a negative perspective, which in turn enhances its overall positive prediction capability. To the best of our knowledge, there are only two concurrent approaches (i.e., CLIPN [66] and LSN [43]) that work in a similar manner: they both enable the negative predictions of CLIP to identify unknown out-of-distribution samples. However, we formulate negative learning from a different and more fundamental perspective, i.e., using negative prediction to directly improve the accuracy of positive prediction.
Contrastive Learning. We also recognize the underlying concept behind our proposed DualAdapter has parallels with existing supervised contrastive learning literature [14, 82, 26]: Our method also aims to maximize similarities between positive image-text/image-image pairs, while simultaneously minimizing similarities between negative pairs during few-shot adaptation. However, our work focuses on a fundamentally different dual learning perspective, i.e., we explicitly train two distinct positive and negative classifiers for ensembling using a single image input at a time, rather than applying supervised contrastive loss to train a single classifier with image pairs as inputs. It is also important to note that contrastive learning is not directly applicable in our fine-tuning setting because we keep the CLIP image encoder frozen. Thereby, the extracted image features are fixed and cannot be pulled together or pushed apart.
Figure 2: An overview of our proposed DualAdapter. We introduce four adapters, each of which builds a cache with positive/negative text/image features that are updated during few-shot training. Given an image to be classified, the classification logit for a specific class increases when the image feature closely aligns with the corresponding features in the positive cache and diverges from those in the negative cache.
3 DualAdapter: Unified Dual-Path Adaptation of VLMs
We introduce a novel framework DualAdapter, as illustrated in Fig. 2, to enable efficient dual-path adaptation of VLMs. Our DualAdapter consists of four adapters: two that positively adapt CLIP to enhance its ability to accurately identify the true class of an input image, and two that negatively adapt CLIP to effectively exclude incorrect candidate classes. Furthermore, we formulate few-shot learning as a noisy problem, and propose an unsupervised similarity-based label refinement approach to select high-quality representative samples to more effectively adapt CLIP to downstream tasks. In the following sections, we first briefly review zero-shot CLIP, then introduce each component of our DualAdapter in detail.
3.1 Preliminary: A Revisit of CLIP
By leveraging contrastive training on billions of image-text pairs from the Internet, CLIP has shown an outstanding ability to align vision and language modalities [48]. In this work, we employ CLIP’s pre-trained visual encoder $\mathcal{F}_{\mathsf{V}}$ and textual encoder $\mathcal{F}_{\mathsf{T}}$ to map the images and textual descriptions into a shared $d$-dimensional embedding space. For a $C$-class classification task, CLIP performs zero-shot predictions by assessing the similarity between the image feature and the text feature specific to each class $c\in\{1,...,C\}$ as follows:
$f_{\mathcal{I}}=\mathcal{F}_{\mathsf{V}}(\mathcal{I}),\quad f_{t_{c}}=\mathcal{F}_{\mathsf{T}}(\mathcal{T}_{c}),\quad\mathbb{P}(y=c|\mathcal{I})=\frac{\exp\left(\mathrm{cos}\left(f_{t_{c}},f_{\mathcal{I}}\right)/t\right)}{\sum\nolimits_{t^{\prime}}{\exp\left(\mathrm{cos}\left(f_{t^{\prime}},f_{\mathcal{I}}\right)/t\right)}},$ (1)
[…]
To streamline this process, a weight matrix can be precomputed and stored in a textual cache, which concatenates the textual features associated with each class, denoted as $\mathsf{T}_{\mathtt{cache}}=[f_{t_{1}}\,f_{t_{2}}\,\cdots\,f_{t_{C}}]^{\top}\in\mathbb{R}^{C\times d}$. Subsequently, we can efficiently obtain the logit $\mathcal{S}$ and the final prediction $\mathbb{P}(y|\mathcal{I})$ via vectorized computation:
$\mathcal{S}=f_{\mathcal{I}}\mathsf{T}_{\mathtt{cache}}^{\quad\,\,\,\,\top}\in\mathbb{R}^{1\times C},\quad\mathbb{P}(y|\mathcal{I})=\mathsf{Softmax}\left(\mathcal{S}\right).$ (2)
3.2 Positive Adaptation
Although the zero-shot CLIP model is trained on web-scale datasets and encodes extensive prior knowledge, it cannot be effectively applied to downstream tasks due to the presence of domain shifts and a deficiency in task-specific knowledge. To address this, we efficiently leverage few-shot annotated samples to distill new, task-specific knowledge. As we discussed in Section 1, our DualAdapter enables few-shot adaptation from both positive and negative perspectives. In this section, we first focus on the positive side, i.e., refining CLIP’s ability to more precisely predict the true class of the input image in a specific downstream task.
Positive Textual Adapter. As shown in Eq. (2), CLIP can make zero-shot predictions utilizing a textual cache, which stores the textual features from positive text descriptions (e.g., “A photo of a {CLASS}”). To avoid ambiguity, we denote this cache as $\mathsf{T}_{\mathtt{cache}}^{+}=[f_{t_{1}}^{+}\,f_{t_{2}}^{+}\,\cdots\,f_{t_{C}}^{+}]^{\top}\in\mathbb{R}^{C\times d}$. With a small set of annotated training images, we can improve this cache-based classifier by enhancing the inter-modal alignment between textual features and the few-shot image features from the same class. Specifically, we follow TaskRes [75] to achieve this by introducing a group of learnable parameters $\mathcal{R}_{\mathsf{T}}^{+}\in\mathbb{R}^{C\times d}$. By incorporating these parameters with the textual cache $\mathsf{T}_{\mathtt{cache}}^{+}$ via element-wise addition in a residual approach, it integrates task-specific knowledge and enhances the inter-modal classifier:
$\mathsf{T}_{\mathtt{cache}}^{+}\leftarrow\mathsf{Normalize}\left(\mathsf{T}_{\mathtt{cache}}^{+}+\mathcal{R}_{\mathsf{T}}^{+}\right),\quad\mathcal{S}_{\mathsf{T}}^{+}=f_{\mathcal{I}}\mathsf{T}_{\mathtt{cache}}^{+\quad\top}\in\mathbb{R}^{1\times C}.$ (3)
Here, $\mathsf{Normalize}$ denotes the L2-normalization applied to each row of the matrix.
Positive Visual Adapter. Inspired by Tip-Adapter [79], we further extend the recognition capability of the CLIP model by constructing a visual cache-based classifier, which operates by measuring intra-modal similarities between the input image feature and few-shot image features. Given a $C$-class $K$-shot training dataset from a specific downstream task, we store all $CK$ image features in a precomputed visual cache, denoted as $\mathsf{V}_{\mathtt{cache}}^{+}=[f_{v_{1}}^{(1)}\,f_{v_{1}}^{(2)}\,\cdots\,f_{v_{C}}^{(K)}]^{\top}\in\mathbb{R}^{CK\times d}$. Their one-hot labels are also correspondingly recorded in $L\in\mathbb{R}^{CK\times C}$. Given an image feature $f_{\mathcal{I}}$ to be classified, we calculate its image-image affinities $\mathcal{A}^{+}$ with all the few-shot training images, then multiplied by their corresponding one-hot labels $L$ to obtain the classification logit $\mathcal{S}_{\mathsf{V}}^{+}$:
$\mathcal{A}^{+}=\exp\left(-\beta\left(1-f_{\mathcal{I}}\mathsf{V}_{\mathtt{cache}}^{+\quad\top}\right)\right)\in\mathbb{R}^{1\times CK},\quad\mathcal{S}_{\mathsf{V}}^{+}=\alpha\mathcal{A}^{+}L\in\mathbb{R}^{1\times C},$ (4)
where $\alpha$ represents a balance factor and $\beta$ denotes a modulating hyper-parameter. Empirically, the value of $\alpha$ should be increased in cases of larger domain shifts, where more domain-specific few-shot knowledge is needed for adaptation.
Symmetrically, we also introduce a set of learnable parameters $\mathcal{R}_{\mathsf{V}}^{+}\in\mathbb{R}^{C\times d}$, which are broadcast to $\mathbb{R}^{CK\times d}$ and added to the positive visual cache:
$\mathsf{V}_{\mathtt{cache}}^{+}\leftarrow\mathsf{Normalize}\left(\mathsf{V}_{\mathtt{cache}}^{+}+\mathcal{R}_{\mathsf{V}}^{+}\right).$ (5)
3.3 Negative Adaptation
Having introduced the positive branch of our DualAdapter, we now explore the negative inference capabilities of VLMs, adapting them to downstream tasks from a negative perspective. Specifically, we craft negative prompts and generate negative image prototypes to direct CLIP towards more confidently excluding incorrect classes based on the given input image.
Negative Textual Adapter. Recall that the positive textual cache, denoted as $\mathsf{T}_{\mathtt{cache}}^{+}$, is constructed based on the positive class descriptor prompt such as “A photo of a {CLASS}.” Conversely, we introduce a set of negative prompts (e.g., “A photo of no {CLASS},” “A photo without {CLASS}”) that convey meanings opposite in semantics [^1] The comprehensive collection of positive and negative prompts utilized for each dataset is available in the Supplementary Materials.. By leveraging the textual features $f_{t}^{-}$ derived from the negative text descriptions linked with these prompts, we again precompute a weight matrix and construct negative textual cache $\mathsf{T}_{\mathtt{cache}}^{-}=[f_{t_{1}}^{-}\,f_{t_{2}}^{-}\,\cdots\,f_{t_{C}}^{-}]^{\top}\in\mathbb{R}^{C\times d}$. Intuitively, if the feature of an input image $f_{\mathcal{I}}$ closely resembles the negative textual feature for a specific class $c$, it strongly suggests that the image does not belong to class $c$, i.e., $\mathbb{P}(y=c|\mathcal{I})\propto 1-f_{\mathcal{I}}^{\top}f_{t_{c}}^{-}$. Considering the zero-shot CLIP model is not designed to perform this kind of negative prediction, we incorporate a learnable residual $\mathcal{R}_{\mathsf{T}}^{-}\in\mathbb{R}^{C\times d}$ to refine the negative text embeddings through supervised task-specific training. The prediction logit is subsequently given by:
$\mathsf{T}_{\mathtt{cache}}^{-}\leftarrow\mathsf{Normalize}\left(\mathsf{T}_{\mathtt{cache}}^{-}+\mathcal{R}_{\mathsf{T}}^{-}\right),\quad\mathcal{S}_{\mathsf{T}}^{-}=\delta_{\mathsf{T}}\left(\mathbf{1}-f_{\mathcal{I}}\mathsf{T}_{\mathtt{cache}}^{-\quad\top}\right)\in\mathbb{R}^{1\times C},$ (6)
where $\delta_{\mathsf{T}}$ is a fixed scaling parameter that adjusts $\mathcal{S}_{\mathsf{T}}^{-}$ to match the mean value of $\mathcal{S}_{\mathsf{T}}^{+}$, and $\mathbf{1}\in\mathbb{R}^{1\times C}$ denotes an all-ones $C$-dimensional vector.
Negative Visual Adapter. In the visual domain, we also similarly pursue some negative image features to enable negative predictions. To achieve this, we naturally treat all $(C-1)K$ images from the other $C-1$ classes as negative samples for a given class $c$. These negative image samples hold the same property as the negative text descriptions: they are explicitly selected to be non-representative of a specific class, which means that the higher the similarity to them, the lower the probability that the image is classified to that class. To mitigate individual biases, we randomly select one image from each of the $C-1$ classes and compute the average of their extracted features to represent the negative image prototypes [56]. In this way, we can get a total of $K$ negative image prototypes for each of the $C$ classes, thereby constructing a negative visual cache $\mathsf{V}_{\mathtt{cache}}^{-}\in\mathbb{R}^{CK\times d}$. Similar to Eq. (4), we compute the affinities between the input image and those negative image prototypes, and calculate the classification logit:
$\mathcal{A}^{-}\!=\!\delta_{\mathsf{V}}\exp\left(-\beta\left(\mathbf{1}-\left(\mathbf{1}-f_{\mathcal{I}}\mathsf{V}_{\mathtt{cache}}^{-\quad\top}\right)\right)\right)$ $\!=\!\delta_{\mathsf{V}}\exp\left(-\beta f_{\mathcal{I}}\mathsf{V}_{\mathtt{cache}}^{-\quad\top}\right)\in\mathbb{R}^{1\times CK},$ (7)
$\mathcal{S}_{\mathsf{V}}^{-}=\alpha\mathcal{A}^{-}L$ $\in\mathbb{R}^{1\times C}.$ (8)
where $\delta_{\mathsf{V}}$ is another fixed scaling parameter that adjusts $\mathcal{A}^{-}$/$\mathcal{S}_{\mathsf{V}}^{-}$ to match the mean value of $\mathcal{A}^{+}$/$\mathcal{S}_{\mathsf{V}}^{+}$, $L\in\mathbb{R}^{CK\times C}$ denotes the corresponding one-hot labels.
To further encode the diverse fine-grained visual knowledge associated with the specific downstream task, we introduce an extra set of learnable residual parameters, $\mathcal{R}_{\mathsf{V}}^{-}\in\mathbb{R}^{C\times d}$, which is similarly broadcast to refine the negative visual cache:
$\mathsf{V}_{\mathtt{cache}}^{-}\leftarrow\mathsf{Normalize}\left(\mathsf{V}_{\mathtt{cache}}^{-}+\mathcal{R}_{\mathsf{V}}^{-}\right).$ (9)
3.4 Unified Dual-Path Adapter
To effectively adapt VLMs to downstream tasks, our DualAdapter combines the functionalities of all four introduced adapters. As illustrated in Figure 3 (Left), we provide an overview of the specific example features stored in the cache for the class dog. By measuring the similarities between the input image feature and these cached example features, our DualAdapter produces a unified final prediction of the probability that the input image represents a dog. The same process is consistently applied across all classes to perform recognition tasks.
DualAdapter Inference. To derive the final classification scores, we aggregate the four outputs from dual-path (i.e., positive and negative) adapters spanning both textual and visual modalities, expressed as:
$\mathcal{S}_{\mathsf{final}}=\lambda\left(\mathcal{S}_{\mathsf{T}}^{+}+\mathcal{S}_{\mathsf{V}}^{+}\right)+(1-\lambda)\left(\mathcal{S}_{\mathsf{T}}^{-}+\mathcal{S}_{\mathsf{V}}^{-}\right).$ (10)
Here, $\lambda$ serves as a tuning hyper-parameter to balance the contribution of positive and negative adapter logits. Throughout the training process, the collection of learnable parameters =$\mathcal{R}=\{\mathcal{R}_{\mathsf{T}}^{+},\mathcal{R}_{\mathsf{V}}^{+},\mathcal{R}_{\mathsf{T}}^{-},\mathcal{R}_{\mathsf{V}}^{-}\}$ is updated via stochastic gradient descent, employing a cross-entropy loss function for optimization.
Figure 3: Visualization of cosine similarities on the validation set of ImageNet [4]. (Left) Example features of the class dog stored in the cache; (Middle) Distribution of pairwise image-text similarities with dual-path features in textual adapters; (Right) Distribution of pairwise image-image similarities with dual-path features in visual adapters.
Effectiveness of Four Adapters. Fig. 3 qualitatively shows the effectiveness of the four adapters we designed. In the middle and right sub-figure, we visualize the distributions of the pairwise cosine similarities between the features of the input image and the textual/visual features stored in the adapters on ImageNet [4] validation set. From Fig. 3, we can observe the following statistical patterns: (1) As we expected, the input image feature is more similar with positive features from the same class (blue $>$ orange) and negative features from other classes (red $>$ green) across both two modalities; (2) Within the textual modality, negative features tend to be less similar to the input image compared to positive features (red $<$ blue, green $<$ orange), since the negative prompts are less common in CLIP [48] training corpus; (3) Within the visual modality, the similarity distribution of negative features occupies an intermediate position between the positive features of the same and different classes (orange $<$ green/red $<$ blue). This is because the negative visual features are constructed by averaging the image features across all other classes, inherently leading to a more generic representation that lacks the distinctiveness characteristic of positive features within a single class.
3.5 Similarity-Based Label Refinement
Previous adapter-style fine-tuning methods [79, 64] generally assume that the few-shot training samples are meticulously curated to provide accurate representation for each class. However, in practical scenarios, this assumption often does not hold true. We recognize that not every image is of high quality or equally representative of its respective class, which can be particularly detrimental in the context of few-shot learning, where the availability of training images is extremely limited. For example, we provide the t-SNE [38] visualization of few-shot image features from the OxfordPets [45] dataset in Fig. 4 (Left). We argue that the few-shot samples are highly noisy: there is even a data point from the Basset Hound class that overlaps with the prototype (i.e., average feature) of Cocker Spaniel.
In response to this, we devise an unsupervised similarity-based label refinement process to address this challenge. Recall that in our visual adapter, we calculate the image-image affinities between the input image and few-shot training images, with all few-shot images being assigned equal weights [^2] Each few-shot image is assigned a one-hot label, corresponding to a weight of 1. as detailed in Eq. (4). As we discussed, this approach is possibly problematic in the noisy few-shot learning setting. Consequently, we extend the visual adapter to assign non-uniform confidence scores according to pairwise similarities, effectively downweighting outliers and emphasizing more representative samples.
Specifically, our refinement process is based on an intuitive assumption: the representative image feature is closer to other image features from the same class than those low-quality outliers [26]. Based on this assumption, given the $K$-shot image features $\{f_{v_{c}}^{(i)}\}_{i=1}^{K}$ from a specific class $c$, we calculate the average cosine similarities of each image feature to others:
$\mathsf{d}_{c}^{(i)}=\frac{1}{K-1}\sum_{j,\,j\neq i}\cos\left(f_{v_{c}}^{(i)},f_{v_{c}}^{(j)}\right)=\frac{1}{K-1}\sum_{j,\,j\neq i}f_{v_{c}}^{(i)\top}f_{v_{c}}^{(j)}.$ (11)
We then assign non-uniform confidence scores $w_{c}^{(i)}$ and compute refined label values $\ell_{c}^{(i)}$ for all $K$-shot image features based on their similarities to others:
$w_{c}^{(i)}=\frac{\exp(\mathsf{d}_{c}^{(i)}/\tau)}{\sum_{i^{\prime}}\exp(\mathsf{d}_{c}^{(i^{\prime})}/\tau)},\quad\ell_{c}^{(i)}=Kw_{c}^{(i)},$ (12)
where $\tau$ is a temperature hyper-parameter to control the intensity of our refinement. In particular, as $\tau\rightarrow\infty$, our refinement becomes inactive, assigning a label value of 1 to all instances. Conversely, as $\tau\rightarrow 0$, our refinement process becomes more selective, assigning the highest label value exclusively to the most representative image feature, while all other features are assigned 0.
This refinement process is applied across each class, wherein the original one-hot labels $L$, are weighted by the new label values computed in Eq. (12), resulting in the updated labels $\mathbb{L}$. This refinement process is incorporated into both the positive and negative visual adapters; therefore we can rewrite Eqs. (4) and (8) as:
$\mathcal{S}_{\mathsf{V}}^{+}=\alpha\mathcal{A}^{+}\mathbb{L}\in\mathbb{R}^{1\times C},\quad\mathcal{S}_{\mathsf{V}}^{-}=\alpha\mathcal{A}^{-}\mathbb{L}\in\mathbb{R}^{1\times C}.$ (13)
Figure 4: Motivation for the proposed similarity-based label refinement. (Left) t-SNE [38] visualization of visual features for 4 representative classes from the OxfordPets [45] dataset, with prototype [56] of each class denoted by the crossmark $\times$; (Middle) Our refined labels for 4 example images from class Beagle; (Right) An illustrative plot of the modeled distributions using one-hot labels and our refined labels for the Beagle class.
Effectiveness of Label Refinement. In Fig. 4 (Middle), we present our refined label values for 4 example images from class Beagle in OxfordPets [45] dataset. Notably, our unsupervised refinement approach functions by measuring the pairwise similarities among 8-shot images from class Beagle, without requiring any extra parameters or additional training. Our similarity-based refinement process allocates higher values to the more representative samples at the top, and lower values to the lower-quality images at the bottom, where the Beagle is either obscured or presented in black and white. In Fig. 4 (Right), we provide an illustration of the modeled visual feature distribution for the Beagle class using one-hot labels and our refined labels. Through our refinement process, we are better equipped to capture the underlying feature distribution (typically, a normal distribution [68, 81]) of the Beagle class, whereas the use of one-hot labels restricts us to modeling merely a uniform distribution.
4 Experiments
In this section, we conduct extensive experiments on two tasks across 15 datasets. These results demonstrate that our DualAdapter outperforms other state-of-the-art methods, exhibiting enhanced adaptation and generalization capabilities.
4.1 Experimental Settings
Tasks and Datasets. To validate the effectiveness of our proposed DualAdapter, we evaluate our proposed method on two standard benchmarking tasks: few-shot learning and domain generalization, respectively. For few-shot learning task, we comprehensively evaluate our method on 11 well-known image classification benchmarks: ImageNet [4], Caltech101 [11], OxfordPets [45], StandfordCars [29], Flowers102 [44], Food101 [1], FGVCAircraft [39], DTD [3], SUN397 [71], EuroSAT [18], and UCF101 [57]. For domain generalization, we investigate the generalization capability of our DualAdapter on 4 variant datasets of ImageNet: ImageNet-V2 [50], ImageNet-Sketch [65], ImageNet-A [20], and ImageNet-R [19].
Implementation Details. Following previous works [83, 79], we adopt ResNet-50 [17] backbone as the visual encoder of CLIP in our experiments by default. We adopt prompt ensembling, leveraging textual prompts from both CLIP [48] and CuPL [46] to enhance model performance. For the negative prompts we used for each dataset, please kindly refer to the Supplementary Materials. We set the hyper-parameters $\lambda$ and $\tau$ as $0.75$ and $1$, respectively. Our DualAdapter is trained using the AdamW [36] optimizer with a cosine scheduler [35]. The batch size is set to 256. For $\mathcal{R}_{\mathsf{T}}^{+}$ and $\mathcal{R}_{\mathsf{V}}^{+}$, the learning rate is set to $0.0001$, while for $\mathcal{R}_{\mathsf{T}}^{-}$ and $\mathcal{R}_{\mathsf{V}}^{-}$, the learning rate is set to $0.0005$. Additionally, our model is trained for 200 epochs on the EuroSAT [18] dataset, and for 20 epochs on all other datasets. To ensure the reliability of our results, we perform each experiment three times using different initialization seeds and report the mean accuracy achieved. All experiments are conducted on a single NVIDIA RTX 6000 Ada GPU.
Baselines. We compare our proposed method with the following state-of-the-art methods: zero-shot and linear probe CLIP [48], CoOp [84], CoCoOp [83], ProGrad [87], CLIP-Adapter [12], Tip-Adapter-F [79], TPT [55], TaskRes [75], and GraphAdapter [32]. For a fair comparison, we directly report the results of these baselines from their respective original papers.
4.2 Few-Shot Learning
In Fig. 5, we compare the few-shot learning performance of our DualAdapter with other state-of-the-art methods on 11 image classification datasets. In the top-left sub-figure, we also present the average classification accuracy across all 11 datasets. The results indicate that our DualAdapter consistently outperforms other methods across various few-shot learning settings by substantial margins. Moreover, our DualAdapter demonstrates more pronounced performance improvements in specialized classification tasks, such as satellite image and texture classification on the EuroSAT [18] and DTD [3] datasets. With 16-shot training on these two datasets, our DualAdapter surpasses Tip-Adapter-F [79] by a notable 3.56% and 3.50%, and enhances the performance of zero-shot CLIP [48] significantly by 45.56% and 30.50%, respectively. For full numerical results, please refer to our Supplementary Materials. Overall, the consistently superior performance on 11 datasets fully demonstrates the effectiveness of our DualAdapter.
Figure 5: Performance comparisons on few-shot learning on 11 image classification datasets. For each dataset, we report the mean accuracy and 95% confidence interval over 3 random seeds of our DualAdapter on 1-/2-/4-/8-/16-shot settings.
4.3 Robustness to Natural Distribution Shifts
Table 1: Performance comparison on robustness to distribution shifts. All the models are trained on 16-shot ImageNet [4] and directed tested on the OOD target datasets. The best results are in bold and the second best are underlined.
Method Source Target
ImageNet -V2 -Sketch -A -R Avg.
Zero-Shot CLIP [48] 60.33 53.27 35.44 21.65 56.00 41.59
Linear Probe CLIP [48] 56.13 45.61 19.13 12.74 34.86 28.09
CoOp [84] 62.95 55.40 34.67 23.06 56.60 42.43
CoCoOp [83] 62.71 55.72 34.48 23.32 57.74 42.82
ProGrad [87] 62.17 54.70 34.40 23.05 56.77 42.23
TPT [55] 60.74 54.70 35.09 26.67 59.11 43.89
TaskRes [75] 64.75 56.47 35.83 22.80 60.70 43.95
GraphAdapter [32] 64.94 56.58 35.89 23.07 60.86 44.10
DualAdapter (Ours) 66.52 57.87 36.38 25.73 61.12 45.28
In Table 1, we compare the generalization performance of our DualAdapter with other state-of-the-art methods in the presence of distribution shifts. Specifically, all the models are trained solely on 16-shot ImageNet [4], and directly tested on 4 out-of-distribution ImageNet variant datasets. As shown in Table 1, our DualAdapter not only achieves state-of-the-art performance on the source dataset but also attains an average performance gain of 1.18% across 4 out-of-distribution (OOD) target datasets. These experimental results demonstrate that by enabling adaptation through both positive and negative paths, our DualAdapter can more effectively prevent overfitting to the source dataset, thereby enhancing robustness against distribution shifts.
4.4 Ablation Studies
[…]
Table 2: Ablation studies for different variants of DualAdapter. We evaluate the few-shot adaptation capabilities of four DualAdapter variants on ImageNet [4].
# Method $\mathcal{R}_{\mathsf{T}}^{+}$ $\mathcal{R}_{\mathsf{T}}^{-}$ $\mathcal{R}_{\mathsf{V}}^{+}$ $\mathcal{R}_{\mathsf{V}}^{-}$ 1-shot 2-shot 4-shot 8-shot 16-shot
[…]
3 ${}^{+}$ 62.83 63.31 63.95 65.13 66.27
4 ${}^{-}$ 62.65 63.07 63.60 64.36 65.12
5 DualAdapter 62.89 63.47 64.12 65.37 66.52
Effects of Different Visual Backbones. We also implement our DualAdapter with various visual encoders, including ResNet [17] and ViT [6], and evaluate its performance against other fine-tuning approaches in Fig. 6 (Left). We can see that our DualAdapter consistently exceeds other methods across all visual backbones, indicating the general effectiveness of our approach.
Figure 6: More ablation results. (Left) Performance comparison of our DualAdapter with other methods on few-shot learning task using different visual backbones; (Middle) Sensitivity analysis of $\lambda$ from Eq. (10) on ImageNet [4]; (Right) Sensitivity analysis of $\tau$ from Eq. (12) on 3 datasets, with the value of $\tau$ presented on a logarithmic scale.
[…]
Scalability to More Shots. When scaling up to more than 16 shots, Tip-Adapter-F [79] limits the cache size to prevent memory overflow. In contrast, our DualAdapter features reduced memory requirements, enabling it to operate without imposing cache size constraints, even in the 128-shot setting. Performance comparisons in Table 4.4 indicate that our DualAdapter outperforms Tip-Adapter-F [79] more significantly in higher shot settings, e.g., by 1.24% in the 128-shot setting.
Efficiency Comparison. We also compare the efficiency of DualAdapter with existing methods in Table 4.4. Our DualAdapter achieves the highest accuracy while also exhibiting advantageous computational efficiency: (1) Compared to prompt-based learning methods such as CoOp [84] and ProGrad [87], our DualAdapter requires approximately 300$\times$ less training time and over 1000$\times$ fewer FLOPs since we do not need to propagate gradients through the textual encoder; (2) Compared to adapter-style fine-tuning methods, our DualAdapter is also competitive in efficiency. It requires 10$\times$ less training time than CLIP-Adapter [12], and demands 3$\times$ fewer FLOPs and 4$\times$ fewer learnable parameters than Tip-Adapter-F [79].
Table 3: Performance comparison with Tip-Adapter(-F) [79] on ImageNet [4] with 16-/32-/64-/128-shot settings.
Table 4: Efficiency comparison with other existing methods on 16-shot ImageNet [4].
Method 16-shot 32-shot 64-shot 128-shot
CLIP [48] ${}^{\dagger}$0-shot: 60.33
Tip-Adapter [79] 62.03 62.51 62.88 63.15
Tip-Adapter-F [79] 65.47 66.58 67.96 69.74
DualAdapter (Ours) 66.52 67.68 69.01 70.98
Method Training Epochs GFLOPs Param. Accuracy
CLIP [48] - - - - 60.33
CoOp [84] 14 hr 200 >10 0.01M 62.95
ProGrad [87] 17 hr 200 >10 0.01M 63.45
CLIP-Adapter [12] 50 min 200 0.004 0.52M 63.59
Tip-Adapter-F [79] 5 min 20 0.030 16.38M 65.51
DualAdapter (Ours) 5 min 20 0.009 4.10M 66.52
Table 4: Efficiency comparison with other existing methods on 16-shot ImageNet [4].
5 Conclusion
In this work, we introduce DualAdapter, a novel approach to effectively adapt vision-language models for downstream tasks by integrating both positive and negative adapters. This innovative approach allows for effective few-shot adaptation from a dual perspective. Furthermore, we develop an unsupervised similarity-based label refinement approach to mitigate the adverse effects of low-quality image samples during the few-shot adaptation process. Comprehensive evaluations on 15 diverse datasets demonstrate that DualAdapter outperforms the state-of-the-art methods in both few-shot learning and domain generalization tasks.
Limitations. We identify two potential limitations of our DualAdapter: (1) Its efficacy in zero-shot situations is constrained due to the scarcity of original negative descriptions in the CLIP training corpus; (2) Similar to other adapter-style approaches such as Tip-Adapter-F [79], our DualAdapter fine-tuned on a specific task cannot be directly applied to another task without additional adaptation. However, recently, Wang et al. [68] shows that adapter-style fine-tuning methods can be extended for these scenarios using the $k$NN algorithm.
Acknowledgement
This work has been funded in part by the Army Research Laboratory (ARL) under grant W911NF-23-2-0007 and W911NF-19-2-0146, and the Air Force Office of Scientific Research (AFOSR) under grants FA9550-18-1-0097 and FA9550-18-1-0251.
Supplementary Material
In this supplementary document, we provide additional details and experimental results to enhance understanding and insights into our proposed DualAdapter. This supplementary document is organized as follows:
  • $\bullet$
    Full numerical results for the few-shot learning task are detailed in Section A.1.
  • $\bullet$
    We compare our DualAdapter with other state-of-the-art methods in domain generalization tasks, utilizing an alternative ViT backbone, in Section A.2.
  • $\bullet$
    More sensitivity analyses of the hyper-parameters are conducted in Section A.3.
  • $\bullet$
    Detailed statistics for all utilized datasets are provided in Section B.1.
  • $\bullet$
    We list the specific positive and negative prompts we used for each dataset in Section B.2.
  • $\bullet$
    Finally, we explore potential avenues for future research in Section C.
A Additional Experimental Results
A.1 Full Numerical Results on Few-Shot Learning
In Fig. 5 in Section 4.2 of the main text, we have evaluated our DualAdapter on few-shot learning task and compared with other state-of-the-art methods. In Table A1, we present the corresponding full numerical results on the few-shot learning task. We also report the 95% confidence interval over 3 random seeds of our DualAdapter to ensure reliability of our results. In the last column, we present the average recognition accuracy over 11 datasets. The results indicate that our DualAdapter consistently outperforms other state-of-the-art methods across various few-shot learning settings by substantial margins.
Our DualAdapter demonstrates superior recognition performance across nearly all tested scenarios, with certain exceptions in lower-shot settings for the Food101 [1] and OxfordPets [45] datasets. We attribute this to the prevalent challenge of overfitting, a common issue not exclusive to our approach but also affecting many existing methods, especially TaskRes [75]: While TaskRes consistently secures the second-best performance across the other 9 datasets, it underperforms significantly on these two. We hypothesize that this issue arises from the noisy training data with intense colors and sometimes wrong labels [1, 45]. However, in this work, we have designed a label refinement mechanism specifically to address this issue. As a result, our DualAdapter secures a relatively robust performance and achieves the second-best on these two datasets.
Table A1: Full numerical results on few-shot learning task. For each dataset, we report the mean accuracy and 95% confidence interval over 3 random seeds of our DualAdapter on 1-/2-/4-/8-/16-shot settings. ${}^{\dagger}$e report the zero-shot performance of CLIP [48] for all settings. For TaskRes [75], we report the results using the enhanced base classifier (i.e., TaskRes*). The best results are in bold and the second are underlined.
Method Setting
Caltech101 [11]
DTD [3]
EuroSAT [18]
FGVCAircraft [39]
Flowers102 [44]
Food101 [1]
ImageNet [4]
OxfordPets [45]
StanfordCars [29]
SUN397 [71]
UCF101 [57]
Avg.
${}^{\dagger}$ero-shot CLIP [48] 1-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
CoOp [84] 87.43 44.13 50.51 9.80 67.90 73.71 57.15 86.51 55.48 60.10 62.10 59.53
CoCoOp [83] 86.01 45.14 35.08 17.81 67.52 77.42 60.84 86.96 57.22 62.28 62.84 59.92
CLIP-Adapter [12] 88.70 45.66 61.51 17.21 73.43 76.77 61.20 85.99 55.14 61.28 62.29 62.65
Tip-Adapter-F [79] 89.38 50.31 59.16 20.83 80.13 77.61 61.32 86.47 58.51 62.51 64.91 64.65
TaskRes [75] 88.80 50.20 61.70 21.41 79.17 74.03 61.90 83.60 59.13 62.33 64.77 64.28
90.87 53.13 67.70 23.61 84.17 77.93 62.89 86.90 61.34 65.13 68.25 67.45
DualAdapter (Ours) (±$\pm 0.33$) (±$\pm 0.48$) (±$\pm 0.47$) (±$\pm 0.18$) (±$\pm 0.34$) (±$\pm 0.22$) (±$\pm 0.13$) (±$\pm 0.31$) (±$\pm 0.28$) (±$\pm 0.08$) (±$\pm 0.23$) (±$\pm 0.28$)
${}^{\dagger}$ero-shot CLIP [48] 2-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
CoOp [84] 87.92 45.04 60.43 18.25 77.47 72.26 55.88 82.36 58.10 59.82 64.13 61.97
CoCoOp [83] 89.47 46.20 38.51 20.22 70.70 78.81 61.86 88.81 58.28 63.50 65.23 61.96
CLIP-Adapter [12] 89.32 51.81 64.11 20.10 81.77 77.20 61.52 86.73 58.71 62.21 67.27 65.52
Tip-Adapter-F [79] 89.81 54.00 65.82 23.47 82.50 77.83 61.69 87.10 62.05 63.55 66.23 66.73
TaskRes [75] 90.27 55.13 65.83 24.13 86.57 75.17 62.63 84.63 63.70 64.97 70.00 67.54
91.19 59.17 74.40 26.22 88.43 78.35 63.47 87.68 64.77 66.73 70.25 70.06
DualAdapter (Ours) (±$\pm 0.26$) (±$\pm 0.31$) (±$\pm 0.87$) (±$\pm 0.37$) (±$\pm 0.18$) (±$\pm 0.17$) (±$\pm 0.06$) (±$\pm 0.28$) (±$\pm 0.21$) (±$\pm 0.33$) (±$\pm 0.48$) (±$\pm 0.31$)
${}^{\dagger}$ero-shot CLIP [48] 4-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
CoOp [84] 89.17 53.38 70.20 21.72 85.81 72.72 59.93 87.22 61.92 63.46 67.08 66.60
CoCoOp [83] 90.31 47.90 63.56 20.56 72.72 79.51 62.52 88.60 59.90 64.90 67.90 65.31
CLIP-Adapter [12] 89.98 57.02 73.18 22.99 87.30 77.93 61.84 87.36 62.26 65.90 68.90 68.61
Tip-Adapter-F [79] 90.67 57.78 73.85 26.01 89.02 78.26 62.52 87.72 64.82 66.13 70.87 69.79
TaskRes [75] 90.97 60.70 73.83 25.70 90.20 76.10 63.57 86.33 67.43 67.27 70.93 70.28
92.21 66.01 76.54 28.95 92.04 78.74 64.12 88.13 67.96 68.59 73.46 72.43
DualAdapter (Ours) (±$\pm 0.21$) (±$\pm 0.49$) (±$\pm 0.59$) (±$\pm 0.29$) (±$\pm 0.25$) (±$\pm 0.11$) (±$\pm 0.17$) (±$\pm 0.26$) (±$\pm 0.36$) (±$\pm 0.29$) (±$\pm 0.51$) (±$\pm 0.32$)
${}^{\dagger}$ero-shot CLIP [48] 8-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
CoOp [84] 90.15 59.88 76.51 25.93 90.84 71.52 60.91 86.40 68.49 65.63 71.81 69.82
CoCoOp [83] 90.14 52.21 64.13 22.03 75.88 79.59 62.40 88.74 60.87 65.37 68.25 66.33
CLIP-Adapter [12] 91.22 60.70 78.34 25.77 91.79 78.01 62.68 87.70 67.78 67.52 73.02 71.32
Tip-Adapter-F [79] 91.54 62.67 77.83 30.21 91.85 78.71 64.00 88.07 69.53 68.80 74.50 72.52
TaskRes [75] 92.40 64.77 79.33 31.48 94.73 76.40 64.67 87.17 71.83 68.73 75.33 73.35
93.40 67.78 81.62 33.90 95.23 79.23 65.37 89.29 72.08 70.93 76.84 75.06
DualAdapter (Ours) (±$\pm 0.23$) (±$\pm 0.55$) (±$\pm 0.54$) (±$\pm 0.63$) (±$\pm 0.26$) (±$\pm 0.21$) (±$\pm 0.09$) (±$\pm 0.21$) (±$\pm 0.51$) (±$\pm 0.15$) (±$\pm 0.39$) (±$\pm 0.34$)
${}^{\dagger}$ero-shot CLIP [48] 16-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
CoOp [84] 91.61 63.11 82.36 31.01 94.39 73.80 62.95 87.30 72.51 69.11 75.70 73.07
CoCoOp [83] 90.90 57.53 70.77 22.40 79.14 79.68 62.71 89.93 62.22 67.21 70.81 68.48
CLIP-Adapter [12] 92.44 66.14 82.76 31.83 93.91 78.21 63.59 87.91 74.12 69.59 76.80 74.30
Tip-Adapter-F [79] 92.93 67.33 83.80 35.50 95.01 79.50 65.51 89.71 75.50 71.31 78.01 75.83
TaskRes [75] 93.43 67.13 84.03 36.30 96.03 77.60 65.73 87.83 76.83 70.67 77.97 75.78
93.77 70.83 87.36 40.27 96.51 79.87 66.52 90.58 77.48 72.32 80.28 77.80
DualAdapter (Ours) (±$\pm 0.33$) (±$\pm 0.77$) (±$\pm 0.84$) (±$\pm 0.53$) (±$\pm 0.58$) (±$\pm 0.34$) (±$\pm 0.13$) (±$\pm 0.38$) (±$\pm 0.68$) (±$\pm 0.24$) (±$\pm 0.26$) (±$\pm 0.50$)
A.2 More Results on Domain Generalization
Table A2: Performance comparison on robustness to distribution shifts. All the models are trained on 16-shot ImageNet [4] and directed tested on the OOD target datasets. The best results are in bold and the second best are underlined.
Method Source Target
ImageNet -V2 -Sketch -A -R Avg.
Zero-Shot CLIP [48] 62.05 54.79 40.82 29.57 65.99 47.79
Linear Probe CLIP [48] 59.58 49.73 28.06 19.67 47.20 36.17
CoOp [84] 66.85 58.08 40.44 30.62 64.45 48.40
TaskRes [75] 68.20 59.20 42.50 31.43 69.33 50.62
GraphAdapter [32] 68.47 59.10 42.70 31.73 69.43 50.74
DualAdapter (Ours) 69.63 59.76 43.41 32.48 69.60 51.31
In Table 1 in Section 4.3 of the main text, we compare the generalization performance of our DualAdapter with other state-of-the-art methods in the presence of distribution shifts with ResNet-50 visual backbone. In Table A2, we present the performance comparison on domain generalization task using ViT-B/32 visual backbone. Consistently, our DualAdapter not only achieves state-of-the-art performance on the source dataset but also attains an average performance gain of 0.57% across 4 out-of-distribution (OOD) target datasets. This verifies that our DualAdapter demonstrates superior generalizability compared to other state-of-the-art methods, independent of the visual backbone utilized.
A.3 More Sensitivity Analyses of Hyper-Parameters
Table A3: Sensitivity of hyper-parameters. All the results are reported on 16-shot ImageNet [4].
$\alpha$ 0.0 0.5 1.0 1.2 1.5 2.0 66.34 66.40 66.47 66.52 66.44 66.36 $\beta$ 1.0 1.5 2.0 2.5 3.0 3.5 66.38 66.44 66.52 66.50 66.48 66.40
Building upon the sensitivity analyses of $\lambda$ and $\tau$ detailed in Section 4.4 of the main text, this section extends our examination to include the sensitivity of parameters $\alpha$ and $\beta$ on 16-shot ImageNet [4]. In our experiments on ImageNet [4], we set the hyper-parameters $\alpha$ and $\beta$ defined in Section 3 to $1.2$ and $2.0$, respectively. To comprehensively investigate the effects of different hyper-parameters, we conducted a sensitivity experiment where we varied each hyper-parameter individually and evaluated the performance on 16-shot ImageNet [4] in Fig. A3. We can see that our choice of $\alpha=1.2$ and $\beta=2.0$ yields the highest performance. Moreover, our DualAdapter maintains robust performance when adjusting these two hyper-parameters, since our DualAdapter includes adapters from both textual and visual modalities and each of our four adapters can work effectively, as we presented in Table 2 in the main text.
B Additional Implementation Details
B.1 Dataset Details
In Table B4, we present the detailed statistics of each dataset we used in our experiments, including the number of classes, the sizes of training, validation and testing sets, and their original tasks.
Table B4: Detailed statistics of datasets used in experiments. Note that the last 4 ImageNet variant datasets are designed for evaluation and only contain the test sets.
Dataset Classes Training Validation Testing Task
Caltech101 [11] 100 4,128 1,649 2,465 Object recognition
DTD [3] 47 2,820 1,128 1,692 Texture recognition
EuroSAT [18] 10 13,500 5,400 8,100 Satellite image recognition
FGVCAircraft [39] 100 3,334 3,333 3,333 Fine-grained aircraft recognition
Flowers102 [44] 102 4,093 1,633 2,463 Fine-grained flowers recognition
Food101 [1] 101 50,500 20,200 30,300 Fine-grained food recognition
ImageNet [4] 1,000 1.28M - 50,000 Object recognition
OxfordPets [45] 37 2,944 736 3,669 Fine-grained pets recognition
StanfordCars [29] 196 6,509 1,635 8,041 Fine-grained car recognition
SUN397 [71] 397 15,880 3,970 19,850 Scene recognition
UCF101 [57] 101 7,639 1,898 3,783 Action recognition
ImageNet-V2 [50] 1,000 - - 10,000 Robustness of collocation
ImageNet-Sketch [65] 1,000 - - 50,889 Robustness of sketch domain
ImageNet-A [20] 200 - - 7,500 Robustness of adversarial attack
ImageNet-R [19] 200 - - 30,000 Robustness of multi-domains
B.2 Positive and Negative Prompts
In Table B5, we detail the specific positive and negative prompts utilized for each dataset. Additionally, as mentioned in Section 4.1, we incorporate prompts from CuPL [46] to further enhance model performance.
Table B5: Positive and negative prompts used in experiments. In addition to these prompts, we also employ CuPL [46] prompts to further enhance performance.
Dataset Positive Prompts Negative Prompts
“itap of a {CLASS}.” “itap without any {CLASS}.”
ImageNet [4] “a bad photo of the {CLASS}.” “a bad photo with no {CLASS} in it.”
ImageNet-V2 [50] “a origami {CLASS}.” “a origami that isn’t a {CLASS}.”
ImageNet-Sketch [65] “a photo of the large {CLASS}.” “a photo with no large {CLASS}.”
ImageNet-A [20] “a {CLASS} in a video game.” “a video game scene without a {CLASS}.”
ImageNet-R [19] “art of the {CLASS}.” “art that doesn’t include a {CLASS}.”
“a photo of the small {CLASS}.” “a photo with no small {CLASS}.”
Caltech101 [11] “a photo of a {CLASS}.” “a photo without {CLASS}.”
DTD [3] “{CLASS} texture.” “not {CLASS} texture.”
EuroSAT [18] “a centered satellite photo of {CLASS}.” “a centered satellite photo without {CLASS}.”
FGVCAircraft [39] “a photo of a {CLASS}, a type of aircraft.” “a photo without {CLASS}, a type of aircraft.”
Flowers102 [44] “a photo of a {CLASS}, a type of flower.” “a photo without {CLASS}, a type of flower.”
Food101 [1] “a photo of {CLASS}, a type of food.” “a photo without {CLASS}, a type of food.”
OxfordPets [45] “a photo of a {CLASS}, a type of pet.” “a photo without {CLASS}, a type of pet.”
StanfordCars [29] “a photo of a {CLASS}.” “a photo of no {CLASS}.”
SUN397 [71] “a photo of a {CLASS}.” “a photo without {CLASS}.”
UCF101 [57] “a photo of a person doing {CLASS}.” “a photo of a person not doing {CLASS}.”
C Future Work
In this work, we introduce the first dual-path adapter-style fine-tuning framework for VLMs. We believe that the similar concept of dual learning can also be applied to prompt-based learning methods, and be extended to fine-tune other foundational models (e.g., other VLMs [24, 74] and LLMs [63, 5]). Besides, we also notice that there are a lot of research efforts dedicated to design better prompts (e.g., using LLM [46, 40, 70]) to fully exploit the capabilities of CLIP. We hope that with our work, future research endeavors can also be directed to investigate the utilization of negative prompts to better activate the negative inference capabilities, further broadening the scope and effectiveness of CLIP.

### v2

@@
-Negative Yields Positive: Unified Dual-Path Adapter for Vision-Language Models
+Enhancing Vision-Language Few-Shot Adaptation with Negative Learning
 Abstract
-Recently, large-scale pre-trained Vision-Language Models (VLMs) have demonstrated great potential in learning open-world visual representations, and exhibit remarkable performance across a wide range of downstream tasks through efficient fine-tuning.
-In this work, we innovatively introduce the concept of dual learning into fine-tuning VLMs, i.e., we not only learn what an image is, but also what an image isn’t.
-Building on this concept, we introduce a novel DualAdapter approach to enable dual-path adaptation of VLMs from both positive and negative perspectives with only limited annotated samples.
-In the inference stage, our DualAdapter performs unified predictions by simultaneously conducting complementary positive selection and negative exclusion across target classes, thereby enhancing the overall recognition accuracy of VLMs in downstream tasks.
-Our extensive experimental results across 15 datasets validate that the proposed DualAdapter outperforms existing state-of-the-art methods on both few-shot learning and domain generalization tasks while achieving competitive computational efficiency.
-Code is available at https://github.com/zhangce01/DualAdapter.
-Keywords:
-Vision-Language Models CLIP Few-Shot Classification Domain Generalization Transfer Learning
+Large-scale pre-trained Vision-Language Models (VLMs) have exhibited impressive zero-shot performance and transferability, allowing them to adapt to downstream tasks in a data-efficient manner.
+However, when only a few labeled samples are available, adapting VLMs to distinguish subtle differences between similar classes in specific downstream tasks remains challenging.
+In this work, we propose a Simple yet effective Negative Learning approach, SimNL, to more efficiently exploit the task-specific knowledge from few-shot labeled samples.
+Unlike previous methods that focus on identifying a set of representative positive features defining “what is a {CLASS}", SimNL discovers a complementary set of negative features that define “what is not a {CLASS}", providing additional insights that supplement the positive features to enhance task-specific recognition capability.
+Further, we identify that current adaptation approaches are particularly vulnerable to potential noise in the few-shot sample set.
+To mitigate this issue, we introduce a plug-and-play few-shot instance reweighting technique to suppress noisy outliers and amplify clean samples for more stable adaptation.
+Our extensive experimental results across 15 datasets validate that the proposed SimNL outperforms existing state-of-the-art methods on both few-shot learning and domain generalization tasks while achieving competitive computational efficiency.
+Code is available at https://github.com/zhangce01/SimNL.
 1 Introduction
-Large-scale pre-trained Vision-Language Models (VLMs), such as CLIP [48] and ALIGN [24], provide a new paradigm for multi-modal learning and open-world visual recognition [78, 8].
-Recently, researchers have demonstrated the robustness and effectiveness of these VLMs in interpreting complex visual and textual inputs, as evidenced by their superior performance in a variety of vision-language tasks, e.g., image captioning [73, 21, 60], visual question answering [52, 25, 31], and other domains [7, 85, 54, 58].
-Current VLMs excel in zero-shot inference [86, 15], but struggle in transferring to specific downstream tasks, primarily due to domain shifts and mismatches between pre-training and task-specific data [10, 42].
-Additionally, the extensive number of parameters and demanding computational costs make it intractable for individual practitioners to re-train or fine-tune the entire models to enhance their task-specific performance [84, 12].
-To address this, recent studies have focused on designing efficient fine-tuning strategies that maintain the core pre-trained encoders frozen while incorporating lightweight adapter [12, 79] modules or optimizing prompts[84, 83].
-These methods enable effective task-specific knowledge extraction from a limited amount of annotated training samples, on top of the inherited foundational knowledge from pre-training.
-Figure 1: Dual-path inference examples.
-We show four sample images from diverse recognition tasks, along with their similarities (scaled by 100) to positive/negative text descriptions after few-shot dual-path adaptation.
-Intuitively, an image is more likely to be classified into a specific class if it exhibits higher positive similarity and lower negative similarity.
-We can observe that while the positive branch alone may struggle to distinguish among some closely related classes, incorporating the negative path, which eliminates certain incorrect classes, enhances its ability to accurately identify the true class.
-In this work, we introduce the innovative concept of dual inference in the context of fine-tuning VLMs.
-Take, for instance, a basic binary pet classification task: distinguishing cats from dogs.
-To identify an image as a cat, we can approach the task from both the positive and negative perspectives: one might directly identify the image as a cat (selection), or alternatively, one can also first determine that it is not a dog (exclusion), thereby indirectly concluding it’s a cat.
-These approaches, namely, positive selection and negative exclusion, complement each other and collectively enhance the overall classification accuracy in general classification tasks.
-Building on this concept, we introduce a novel DualAdapter approach for few-shot adaptation of VLMs that conducts positive selection and negative exclusion across a set of target classes.
-While CLIP typically performs classification tasks using prompts such as “A photo of a {CLASS}”, we recognize its ability to understand negation, employing prompts like “A photo of no {CLASS}”.
-Utilizing this capability, we design our DualAdapter to adapt VLMs from a dual perspective: we propose to learn not only what an image is but also what an image isn’t by matching the input image feature with the positive and negative text features.
-Intuitively, our approach first determines a set of likely candidate classes through image classification but, in a second step, refines these classification logits by asking the inverse question of what an image isn’t across the entire set of target classes.
-In Fig.
-1, we showcase several dual-path inference examples to illustrate that when positive selection struggles with distinguishing between similar classes, employing negative exclusion can provide further insights to improve recognition accuracy.
-Moreover, in DualAdapter, we symmetrically apply dual-path adaptation to the visual modality, treating image features of a specific class as positive and those from all other classes as negative.
-Recognizing the variability in image quality and representativeness, where some image features are noisy and not equally representative of their respective classes, we implement an unsupervised similarity-based label refinement technique.
-This method assigns confidence scores to each image, reducing the impact of outliers or less representative samples during few-shot adaptation, ensuring more accurate and reliable classification.
-To empirically validate DualAdapter, we perform thorough evaluations on few-shot learning and domain generalization tasks using 15 diverse recognition datasets.
-These experimental results demonstrate the effectiveness of DualAdapter in adapting VLMs for downstream tasks and verify its superior robustness to distribution shifts.
-Additionally, our efficiency benchmarks indicate that DualAdapter achieves competitive efficiency compared to other state-of-the-art methods.
-In summary, our primary contributions are listed as follows:
-  • $\bullet$
-    We explore and exploit the negative inference capabilities of VLMs, and introduce a unique dual-path adaptation approach (i.e., DualAdapter) for CLIP.
-  • $\bullet$
-    To mitigate the impact of noisy samples in few-shot adaptation, we introduce a novel unsupervised similarity-based label refinement technique.
-  • $\bullet$
-    Through extensive experiments, we demonstrate that our DualAdapter outperforms other state-of-the-art methods in both adaptation performance and generalizability across 15 diverse recognition datasets.
+Over the past decade, deep learning models have achieved remarkable progress in various vision tasks [41, 42], largely due to training on extensive supervised datasets [6, 27].
+However, in real-world scenarios, acquiring such large-scale datasets in specific domains (e.g., satellite and aircraft images) is often impractical due to the time-consuming and costly nature of the collection and annotation process.
+As a result, there is often not enough data in these domains to capture the variability of the novel classes involved, making it challenging to develop robust models from scratch that can effectively generalize to unseen data [58, 47].
+Recent advances in Vision-Language Models (VLMs), such as CLIP [38] and ALIGN [21], provide a promising alternative approach for building robust models in a low-data regime.
+Specifically, by utilizing large-scale pre-training on web-scale datasets, these VLMs have demonstrated remarkable zero-shot performance and transferability [38, 64, 67, 26], which enables the adaptation of these models to downstream tasks in a data-efficient manner.
+Currently, researchers have developed two primary few-shot adaptation strategies for VLMs: prompt-based learning [74, 73, 45, 50, 72] and adapter-style fine-tuning [10, 71, 25, 65, 70], to enable effective task-specific knowledge extraction.
+However, due to the very limited number of annotated samples available, these methods still struggle to discern subtle differences between similar classes in specific downstream tasks.
+Figure 1: Negative learning provides complementary information for more accurate recognition.
+(Left) Grad-CAM [44] visualization of the similarity heatmaps with the learned positive and negative features of the ground-truth class; (Middle) Similarities (scaled by 100) to the learned positive and negative features of five similar candidate classes.
+While the positive branch alone may fail to distinguish among some closely related classes, incorporating the negative classifier, which eliminates certain incorrect classes, enhances the model’s ability to accurately identify the true class; (Right) Performance comparisons with other state-of-the-art methods in 16-shot scenarios.
+In this work, inspired by the negative learning literature [20, 66, 49], we propose a simple yet effective negative learning approach, SimNL, for more effectively adapting VLMs to downstream tasks.
+Specifically, while previous adaptation methods [74, 73, 71] typically focus on identifying a set of representative features that define “what is a {CLASS}" during few-shot adaptation, we propose to discover a complementary set of negative features that define “what is not a {CLASS}" to better exploit the limited task-specific knowledge from the few-shot samples.
+As shown in Figure 1 (Left), this complementary set of negative features guides our model to pay attention to more diverse attributes when classifying an image, evaluating both the presence of class-specific characteristics and the absence of characteristics not associated with the class.
+Building on these discovered features, our SimNL framework incorporates two coordinated classifiers: one performing similarity matching with the positive features and the other performing dissimilarity matching with the negative features.
+In Figure 1 (Middle), we illustrate that when the positive classifier struggles with distinguishing between similar classes, employing the negative one can provide further insights to improve recognition accuracy.
+The performance comparisons in Figure 1 (Right) further highlight that our SimNL method, by incorporating an additional negative classifier, consistently achieves superior few-shot adaptation performance across 11 various datasets when compared to existing state-of-the-art approaches.
+Another important challenge often overlooked in existing adaptation methods is the presence of noise in the few-shot sample set.
+Most existing approaches assume that each few-shot sample is carefully curated to accurately represent its class, but such assumptions rarely holds in practice.
+Given the extremely small size of the few-shot sample set, the models trained from these samples are particularly vulnerable to noise.
+To address this important issue, we extend SimNL by reweighting each few-shot instance with a confidence score, which suppresses outliers and amplifies clean samples to ensure stable adaptation.
+This plug-and-play technique can also be applied to other adapter-style fine-tuning methods, such as Tip-Adapter-F [71], making them more robust to noisy data.
+To empirically validate the effectiveness of our SimNL, we conduct comprehensive evaluations on few-shot learning and domain generalization tasks across 15 diverse recognition datasets.
+These experimental results demonstrate the effectiveness of SimNL in adapting VLMs for downstream tasks and verify its superior robustness to distribution shifts.
+We also demonstrate that our proposed instance reweighting significantly enhances the model’s robustness to label noises in the few-shot sample set.
+Additionally, SimNL also exhibits competitive efficiency.
+Our primary contributions are summarized as follows:
+  • We propose a simple yet effective negative learning approach, i.e., SimNL, to efficiently adapt CLIP to downstream tasks.
+Specifically, SimNL introduces an innovative application of negative learning to adapter-style fine-tuning for few-shot adaptation of VLMs.
+  • To mitigate the impact of noisy samples in few-shot adaptation, we introduce a plug-and-play few-shot instance reweighting technique that assigns non-uniform confidence scores to each sample.
+  • Through extensive experiments, we demonstrate that our SimNL consistently outperforms other state-of-the-art methods across 15 diverse recognition datasets.
 2 Related Work
-Vision-Language Models.
+Efficient Adaptation for VLMs.
 In recent years, extensive Vision-Language Models (VLMs) have been developed to bridge the vision and language modalities through large-scale pre-training [38, 21].
-Generally, these pre-trained VLMs can be classified into four categories based on their pre-training objectives: masked language modeling [37, 30, 27], masked region prediction [37, 30, 59], image-text matching [30, 27, 61, 31], and contrastive learning [48, 24, 74, 77, 33].
-Researchers have demonstrated that these pre-trained VLMs exhibit remarkable open-world visual understanding capabilities [8], achieving state-of-the-art performance across a range of open-world tasks, such as zero-shot learning [48, 51, 41], open-world segmentation [34, 72, 49], and open-world detection [53, 16, 9].
-In this work, we mainly focus on contrastive-based VLMs, especially CLIP [48], which has become the current mainstream approach in this field.
-Nonetheless, we believe that the core concept of this work is adaptable and can be extended to other VLMs as well.
-Efficient Fine-Tuning for VLMs.
 Given the substantial size of the VLMs, recent research efforts [74, 10, 71, 69] are focusing on the development of lightweight fine-tuning techniques to efficiently adapt VLMs for downstream visual tasks.
-These methods can generally be divided into two primary categories: prompt-based learning and adapter-style fine-tuning.
-Specifically, prompt-based learning methods [84, 83, 2, 62] aim to optimize the input prompts from downstream data, while adapter-style fine-tuning approaches [12, 79, 64, 75, 88] directly tune the extracted visual and textual representations.
-Our proposed DualAdapter follows the line of adapter-style fine-tuning, but it distinguishes itself from previous work [79, 75] by incorporating negative adapters for the first time.
-Moreover, we empirically demonstrate that leveraging negative cues from CLIP can effectively improve both few-shot classification performance and generalization capability.
+These methods can generally be divided into two categories: prompt-based learning and adapter-style fine-tuning.
+Specifically, prompt-based learning methods [74, 73, 4, 50, 61] aim to optimize the input prompts from downstream data, while adapter-style fine-tuning approaches [10, 71, 52, 65, 76] directly tune the extracted visual and textual representations.
+In this work, we aim to enhance adapter-style fine-tuning by incorporating negative learning into vision-language few-shot adaptation.
+We also empirically validate that leveraging negative cues from CLIP can effectively improve both few-shot classification performance and generalization capability.
+Few-Shot Learning.
+Few-shot learning aims to quickly adapt a model to new categories using only a few examples.
+Traditional few-shot learning methods primarily rely on meta-learning and can be roughly separated into two groups: metric-based methods [54, 46, 68] and optimization-based methods [39, 43, 12].
+However, these methods depend on training with base datasets, limiting their applicability in real-world scenarios.
+Recent developments in large-scale pre-trained VLMs [38, 1] present a promising alternative due to their exceptional zero-shot capabilities.
+Researchers have demonstrated that with efficient adaptation, VLMs can also excel in few-shot learning tasks [74, 73, 71, 17].
 Negative Learning.
 Given that obtaining typical positive labels (which indicate the categories an image belongs to) can be costly and labor-intensive, researchers have proposed an alternative approach known as the indirect negative learning [20, 66] paradigm.
-This method focuses on learning from easily accessible negative/complementary labels, which specify the categories to which an image does not belong.
+This method focuses on learning from negative/complementary labels, which specify the categories to which an image does not belong.
 In recent years, negative learning has been effectively applied to various vision applications, such as image recognition [20, 23, 11], few-shot learning [18, 60], and semantic segmentation [57, 37].
-In this work, empowered by the strong image-text association capabilities of the advanced pre-trained CLIP [48] model, we enable negative learning within CLIP without the need for explicit negative labels.
-Instead, we design negative prompts and negative image samples to guide the CLIP model to predict from a negative perspective, which in turn enhances its overall positive prediction capability.
-To the best of our knowledge, there are only two concurrent approaches (i.e., CLIPN [66] and LSN [43]) that work in a similar manner: they both enable the negative predictions of CLIP to identify unknown out-of-distribution samples.
-However, we formulate negative learning from a different and more fundamental perspective, i.e., using negative prediction to directly improve the accuracy of positive prediction.
-Contrastive Learning.
-We also recognize the underlying concept behind our proposed DualAdapter has parallels with existing supervised contrastive learning literature [14, 82, 26]: Our method also aims to maximize similarities between positive image-text/image-image pairs, while simultaneously minimizing similarities between negative pairs during few-shot adaptation.
-However, our work focuses on a fundamentally different dual learning perspective, i.e., we explicitly train two distinct positive and negative classifiers for ensembling using a single image input at a time, rather than applying supervised contrastive loss to train a single classifier with image pairs as inputs.
-It is also important to note that contrastive learning is not directly applicable in our fine-tuning setting because we keep the CLIP image encoder frozen.
-Thereby, the extracted image features are fixed and cannot be pulled together or pushed apart.
-Figure 2: An overview of our proposed DualAdapter.
-We introduce four adapters, each of which builds a cache with positive/negative text/image features that are updated during few-shot training.
-Given an image to be classified, the classification logit for a specific class increases when the image feature closely aligns with the corresponding features in the positive cache and diverges from those in the negative cache.
-3 DualAdapter: Unified Dual-Path Adaptation of VLMs
-We introduce a novel framework DualAdapter, as illustrated in Fig.
-2, to enable efficient dual-path adaptation of VLMs.
-Our DualAdapter consists of four adapters: two that positively adapt CLIP to enhance its ability to accurately identify the true class of an input image, and two that negatively adapt CLIP to effectively exclude incorrect candidate classes.
-Furthermore, we formulate few-shot learning as a noisy problem, and propose an unsupervised similarity-based label refinement approach to select high-quality representative samples to more effectively adapt CLIP to downstream tasks.
-In the following sections, we first briefly review zero-shot CLIP, then introduce each component of our DualAdapter in detail.
+In this work, empowered by the strong negative understanding capabilities of the large-scale pre-trained CLIP model [56, 32, 19], we enable an alternative approach to negative learning, where we explicitly train a negative classifier without the need for complementary labels.
+We provide more discussions on the differences between our method and traditional negative learning in Section B.1.
+We also note that recent work on TDA [22] shares a similar approach to ours, as both methods incorporate negative caches to enhance the generalizability of VLMs.
+However, their method focuses on test-time adaptation, where training data is unavailable, making it unsuitable for direct application in few-shot adaptation scenarios.
+3 Method
+In this section, we introduce SimNL, a novel approach to vision-language few-shot adaptation, as shown in Figure 2.
+Figure 2: An overview of our proposed SimNL.
+We construct and learn the positive and negative CLIP-based classifiers across visual and textual modalities.
+Given an image to be classified, the classification logit for a specific class increases when the image feature $f_{v}$ closely aligns with the corresponding positive features $f_{t}^{+},f_{v}^{+}$ and diverges from negative features $f_{t}^{-},f_{v}^{-}$.
+3.1 Problem Formulation
+We consider the problem of $C$-way $K$-shot few-shot classification problem, in which only $K$ labeled examples are provided for each of the $C$ classes.
+Let the feature space be $\mathcal{X}$ and the label space be $\mathcal{Y}=\{1,...,C\}$, the few-shot instance $\boldsymbol{x}\in\mathcal{X}$ and its labels $y\in\mathcal{Y}$ are sampled from probability distribution $P(\boldsymbol{x},y)$.
+The corresponding one-hot encoded labels are represented by $\boldsymbol{y}\in\{0,1\}^{C}$.
+Positive Learning.
+In this study, we employ a CLIP-based classifier with parameters $\theta$, denoted as $\mathcal{F}_{\theta}:\mathcal{X}\rightarrow\mathbb{R}^{C}$, to project each input into a $C$-dimensional score space.
+The classification probabilities are obtained using $\boldsymbol{p}=\texttt{Softmax}\left(\mathcal{F}_{\theta}(\boldsymbol{x})\right)\in\Delta^{C-1}$, where $\Delta^{C-1}$ represents the $(C-1)$-dimensional probability simplex.
+Our goal is to learn a robust classifier $\mathcal{F}_{\theta}$ that minimizes the expected risk; that is,
+$\min_{\theta}\mathcal{R}(\mathcal{F}_{\theta})=\mathbb{E}_{(\boldsymbol{x},y)\sim P(\boldsymbol{x},y)}\left[\mathcal{L}\left(\mathcal{F}_{\theta}(\boldsymbol{x}),y\right)\right],$
+$\text{where}\,\,\mathcal{L}\left(\mathcal{F}_{\theta}(\boldsymbol{x}),y\right)=-\sum_{k=1}^{C}\boldsymbol{y}_{k}\log\boldsymbol{p}_{k},$ (1)
+where $\boldsymbol{p}_{k}$ denotes the probability for the $k$-th class and $\mathcal{L}\left(\mathcal{F}_{\theta}(\boldsymbol{x}),y\right)$ computes the cross-entropy loss.
+We aim to maximize the probability $\boldsymbol{p}_{y}\rightarrow 1$ for the true label.
+Negative Learning.
+In this work, we apply the concept of negative learning to vision-language few-shot adaptation by employing a distinct CLIP-based negative classifier $\mathcal{G}_{\varphi}:\mathcal{X}\rightarrow\mathbb{R}^{C}$ with parameters $\varphi$.
+This classifier predicts the negative probability $\bar{\boldsymbol{p}}=\texttt{Softmax}(\mathcal{G}_{\varphi}(\boldsymbol{x}))$ that the image does not belong to specific classes.
+The expected classification risk for $\mathcal{G}_{\varphi}$ can be written as
+$\min_{\varphi}\mathcal{R}(\mathcal{G}_{\varphi})=\mathbb{E}_{(\boldsymbol{x},y)\sim P(\boldsymbol{x},y)}\left[\mathcal{L}\left(\mathcal{G}_{\varphi}(\boldsymbol{x}),y\right)\right],$
+$\text{where}\,\,\mathcal{L}\left(\mathcal{G}_{\varphi}(\boldsymbol{x}),y\right)=-\sum_{k=1}^{C}\boldsymbol{y}_{k}\log(1-\bar{\boldsymbol{p}}_{k}).$ (2)
+By optimizing this risk, we aim to reduce the negative probability $\bar{\boldsymbol{p}}_{y}\rightarrow 0$ for the true label.
+Finally, we ensemble the optimized classifiers from both positive and negative learning to create an enhanced classifier $\mathcal{F}_{\theta}\oplus\mathcal{G}_{\varphi}:\mathcal{X}\rightarrow\mathbb{R}^{C}$ for testing.
+This is equivalent to combining the positive and negative predicted probabilities as $\hat{\boldsymbol{p}}=\lambda\boldsymbol{p}+(1-\lambda)(1-\bar{\boldsymbol{p}})$, where $\lambda$ is a balancing hyper-parameter.
+Differences with Contrastive Learning.
+Note that the concept of “negative" differs in our negative learning approach compared to contrastive learning: (1) In our negative learning, “negative" specifically refers to a negative classifier.
+We explicitly train another negative classifier $\mathcal{G}_{\varphi}$ to ensemble with the positive classifier; (2) In contrastive learning, “negative" refers to the negative sample pairs that are constructed and utilized during training.
+In Appendix B.2, we show that the training objectives for both the positive and negative CLIP-based classifiers are contrastive in nature.
 3.2 Preliminary: A Revisit of CLIP
-By leveraging contrastive training on billions of image-text pairs from the Internet, CLIP has shown an outstanding ability to align vision and language modalities [48].
-In this work, we employ CLIP’s pre-trained visual encoder $\mathcal{F}_{\mathsf{V}}$ and textual encoder $\mathcal{F}_{\mathsf{T}}$ to map the images and textual descriptions into a shared $d$-dimensional embedding space.
-For a $C$-class classification task, CLIP performs zero-shot predictions by assessing the similarity between the image feature and the text feature specific to each class $c\in\{1,...,C\}$ as follows:
-$f_{\mathcal{I}}=\mathcal{F}_{\mathsf{V}}(\mathcal{I}),\quad f_{t_{c}}=\mathcal{F}_{\mathsf{T}}(\mathcal{T}_{c}),\quad\mathbb{P}(y=c|\mathcal{I})=\frac{\exp\left(\mathrm{cos}\left(f_{t_{c}},f_{\mathcal{I}}\right)/t\right)}{\sum\nolimits_{t^{\prime}}{\exp\left(\mathrm{cos}\left(f_{t^{\prime}},f_{\mathcal{I}}\right)/t\right)}},$ (1)
-[…]
-To streamline this process, a weight matrix can be precomputed and stored in a textual cache, which concatenates the textual features associated with each class, denoted as $\mathsf{T}_{\mathtt{cache}}=[f_{t_{1}}\,f_{t_{2}}\,\cdots\,f_{t_{C}}]^{\top}\in\mathbb{R}^{C\times d}$.
-Subsequently, we can efficiently obtain the logit $\mathcal{S}$ and the final prediction $\mathbb{P}(y|\mathcal{I})$ via vectorized computation:
-$\mathcal{S}=f_{\mathcal{I}}\mathsf{T}_{\mathtt{cache}}^{\quad\,\,\,\,\top}\in\mathbb{R}^{1\times C},\quad\mathbb{P}(y|\mathcal{I})=\mathsf{Softmax}\left(\mathcal{S}\right).$ (2)
-3.2 Positive Adaptation
-Although the zero-shot CLIP model is trained on web-scale datasets and encodes extensive prior knowledge, it cannot be effectively applied to downstream tasks due to the presence of domain shifts and a deficiency in task-specific knowledge.
-To address this, we efficiently leverage few-shot annotated samples to distill new, task-specific knowledge.
-As we discussed in Section 1, our DualAdapter enables few-shot adaptation from both positive and negative perspectives.
-In this section, we first focus on the positive side, i.e., refining CLIP’s ability to more precisely predict the true class of the input image in a specific downstream task.
-Positive Textual Adapter.
+In this work, our classifiers are based on CLIP’s pre-trained visual encoder $\mathcal{F}_{\mathsf{V}}$ and textual encoder $\mathcal{F}_{\mathsf{T}}$.
+Specifically, CLIP [38] performs zero-shot predictions by assessing the similarity between the image feature and the text feature specific to each class $c\in\{1,...,C\}$ as follows:
+$f_{v}=\mathcal{F}_{\mathsf{V}}(\boldsymbol{x}),\quad f_{t_{c}}=\mathcal{F}_{\mathsf{T}}(\mathcal{T}_{c}),$
+$\mathbb{P}(y=c|\boldsymbol{x})=\frac{\exp\left(\mathrm{cos}\left(f_{t_{c}},f_{v}\right)/t\right)}{\sum\nolimits_{t^{\prime}}{\exp\left(\mathrm{cos}\left(f_{t^{\prime}},f_{v}\right)/t\right)}},$ (3)
+where $\boldsymbol{x}$ is the input image, and $\mathcal{T}_{c}$ represents the text description for class $c$ (e.g., “A photo of a {CLASSc}”).
+The parameter $t$ is the temperature hyperparameter, and $\mathrm{cos}\left(f_{t},f_{v}\right)=f_{v}^{\top}f_{t}$ computes the cosine similarity.
+To streamline this process, we can precompute a textual cache, which concatenates the textual features associated with each class, denoted as $\mathsf{T}_{\mathtt{cache}}=[f_{t_{1}}\,f_{t_{2}}\,\cdots\,f_{t_{C}}]^{\top}\in\mathbb{R}^{C\times d}$.
+Subsequently, we can obtain the final prediction $\mathbb{P}(y|\boldsymbol{x})$ via vectorized similarity matching:
+$\mathcal{S}=f_{v}\mathsf{T}_{\mathtt{cache}}^{\quad\,\,\,\,\top}\in\mathbb{R}^{C},\quad\mathbb{P}(y|\boldsymbol{x})=\mathtt{Softmax}\left(\mathcal{S}\right).$ (4)
+3.3 Learning the Positive Classifier
+In this section, we focus on the positive perspective, namely, constructing and optimizing a learnable CLIP-based positive classifier.
+This similarity-based classifier is designed to identify a set of representative features that enhance the accuracy of predicting the true class of the input image.
+Positive Textual Branch.
 As shown in Eq.
@@
 To avoid ambiguity, we denote this cache as $\mathsf{T}_{\mathtt{cache}}^{+}=[f_{t_{1}}^{+}\,f_{t_{2}}^{+}\,\cdots\,f_{t_{C}}^{+}]^{\top}\in\mathbb{R}^{C\times d}$.
-With a small set of annotated training images, we can improve this cache-based classifier by enhancing the inter-modal alignment between textual features and the few-shot image features from the same class.
-Specifically, we follow TaskRes [75] to achieve this by introducing a group of learnable parameters $\mathcal{R}_{\mathsf{T}}^{+}\in\mathbb{R}^{C\times d}$.
-By incorporating these parameters with the textual cache $\mathsf{T}_{\mathtt{cache}}^{+}$ via element-wise addition in a residual approach, it integrates task-specific knowledge and enhances the inter-modal classifier:
-$\mathsf{T}_{\mathtt{cache}}^{+}\leftarrow\mathsf{Normalize}\left(\mathsf{T}_{\mathtt{cache}}^{+}+\mathcal{R}_{\mathsf{T}}^{+}\right),\quad\mathcal{S}_{\mathsf{T}}^{+}=f_{\mathcal{I}}\mathsf{T}_{\mathtt{cache}}^{+\quad\top}\in\mathbb{R}^{1\times C}.$ (3)
-Here, $\mathsf{Normalize}$ denotes the L2-normalization applied to each row of the matrix.
-Positive Visual Adapter.
-Inspired by Tip-Adapter [79], we further extend the recognition capability of the CLIP model by constructing a visual cache-based classifier, which operates by measuring intra-modal similarities between the input image feature and few-shot image features.
-Given a $C$-class $K$-shot training dataset from a specific downstream task, we store all $CK$ image features in a precomputed visual cache, denoted as $\mathsf{V}_{\mathtt{cache}}^{+}=[f_{v_{1}}^{(1)}\,f_{v_{1}}^{(2)}\,\cdots\,f_{v_{C}}^{(K)}]^{\top}\in\mathbb{R}^{CK\times d}$.
+With a small set of annotated training images, we can update the textual features by introducing a group of learnable residual parameters $\mathcal{R}_{\mathsf{T}}^{+}\in\mathbb{R}^{C\times d}$ to integrate task-specific knowledge:
+$\mathsf{T}_{\mathtt{cache}}^{+}\leftarrow\mathtt{Normalize}\left(\mathsf{T}_{\mathtt{cache}}^{+}+\mathcal{R}_{\mathsf{T}}^{+}\right),$
+$\mathcal{S}_{\mathsf{T}}^{+}=f_{v}\mathsf{T}_{\mathtt{cache}}^{+\quad\top}\in\mathbb{R}^{C}.$ (5)
+Here, $\mathtt{Normalize}$ denotes the L2-normalization applied to each row of the matrix.
+Positive Visual Branch.
+We further extend the recognition capability of the CLIP model by constructing a visual cache-based classifier, which operates by measuring image-image similarities between the input image feature and few-shot image features.
+Specifically, we store all $CK$ image features in a precomputed visual cache, denoted as $\mathsf{V}_{\mathtt{cache}}^{+}=[f_{v_{1}}^{(1)}\,f_{v_{1}}^{(2)}\,\cdots\,f_{v_{C}}^{(K)}]^{\top}\in\mathbb{R}^{CK\times d}$.
 Their one-hot labels are also correspondingly recorded in $L\in\mathbb{R}^{CK\times C}$.
-Given an image feature $f_{\mathcal{I}}$ to be classified, we calculate its image-image affinities $\mathcal{A}^{+}$ with all the few-shot training images, then multiplied by their corresponding one-hot labels $L$ to obtain the classification logit $\mathcal{S}_{\mathsf{V}}^{+}$:
-$\mathcal{A}^{+}=\exp\left(-\beta\left(1-f_{\mathcal{I}}\mathsf{V}_{\mathtt{cache}}^{+\quad\top}\right)\right)\in\mathbb{R}^{1\times CK},\quad\mathcal{S}_{\mathsf{V}}^{+}=\alpha\mathcal{A}^{+}L\in\mathbb{R}^{1\times C},$ (4)
-where $\alpha$ represents a balance factor and $\beta$ denotes a modulating hyper-parameter.
-Empirically, the value of $\alpha$ should be increased in cases of larger domain shifts, where more domain-specific few-shot knowledge is needed for adaptation.
-Symmetrically, we also introduce a set of learnable parameters $\mathcal{R}_{\mathsf{V}}^{+}\in\mathbb{R}^{C\times d}$, which are broadcast to $\mathbb{R}^{CK\times d}$ and added to the positive visual cache:
-$\mathsf{V}_{\mathtt{cache}}^{+}\leftarrow\mathsf{Normalize}\left(\mathsf{V}_{\mathtt{cache}}^{+}+\mathcal{R}_{\mathsf{V}}^{+}\right).$ (5)
-3.3 Negative Adaptation
-Having introduced the positive branch of our DualAdapter, we now explore the negative inference capabilities of VLMs, adapting them to downstream tasks from a negative perspective.
-Specifically, we craft negative prompts and generate negative image prototypes to direct CLIP towards more confidently excluding incorrect classes based on the given input image.
-Negative Textual Adapter.
-Recall that the positive textual cache, denoted as $\mathsf{T}_{\mathtt{cache}}^{+}$, is constructed based on the positive class descriptor prompt such as “A photo of a {CLASS}.” Conversely, we introduce a set of negative prompts (e.g., “A photo of no {CLASS},” “A photo without {CLASS}”) that convey meanings opposite in semantics [^1] The comprehensive collection of positive and negative prompts utilized for each dataset is available in the Supplementary Materials..
+Given an image feature $f_{v}$ to be classified, we calculate its image-image affinities and obtain the logits $\mathcal{S}_{\mathsf{V}}^{+}$:
+$\mathsf{V}_{\mathtt{cache}}^{+}\leftarrow\mathsf{Normalize}\left(\mathsf{V}_{\mathtt{cache}}^{+}+\mathcal{R}_{\mathsf{V}}^{+}\right),$
+$\mathcal{S}_{\mathsf{V}}^{+}=\mathcal{A}\left(f_{v}\mathsf{V}_{\mathtt{cache}}^{+\quad\top}\right)L\in\mathbb{R}^{C},$ (6)
+where $\mathcal{A}\left(f_{v}\mathsf{V}_{\mathtt{cache}}^{+\quad\top}\right)=\alpha\exp\left(-\beta\left(1-f_{v}\mathsf{V}_{\mathtt{cache}}^{+\quad\top}\right)\right)$ calculates the affinity, $\alpha$ represents a balance scalar and $\beta$ denotes a modulating parameter.
+Learnable residue $\mathcal{R}_{\mathsf{V}}^{+}\in\mathbb{R}^{C\times d}$ is also introduced, which is broadcast to $\mathbb{R}^{CK\times d}$ and added to the visual cache to refine the visual features.
+3.4 Learning the Negative Classifier
+Having introduced the positive classifier, we now explore constructing and learning a negative dissimilarity-based classifier, which directs CLIP towards more confidently excluding incorrect classes based on the given input image.
+Conceptually, our goal is to mine a general negative feature for each class $c$, which is absent in samples from class $c$ but present in samples from all other classes.
+Negative Textual Branch.
+Recall that the positive textual cache, denoted as $\mathsf{T}_{\mathtt{cache}}^{+}$, is constructed based on the positive class descriptor prompt such as “A photo of a {CLASS}.” Conversely, we introduce a set of negative prompts (e.g., “A photo of not a {CLASS},” “A photo without {CLASS}”) that convey opposite semantics meanings.
 By leveraging the textual features $f_{t}^{-}$ derived from the negative text descriptions linked with these prompts, we again precompute a weight matrix and construct negative textual cache $\mathsf{T}_{\mathtt{cache}}^{-}=[f_{t_{1}}^{-}\,f_{t_{2}}^{-}\,\cdots\,f_{t_{C}}^{-}]^{\top}\in\mathbb{R}^{C\times d}$.
-Intuitively, if the feature of an input image $f_{\mathcal{I}}$ closely resembles the negative textual feature for a specific class $c$, it strongly suggests that the image does not belong to class $c$, i.e., $\mathbb{P}(y=c|\mathcal{I})\propto 1-f_{\mathcal{I}}^{\top}f_{t_{c}}^{-}$.
-Considering the zero-shot CLIP model is not designed to perform this kind of negative prediction, we incorporate a learnable residual $\mathcal{R}_{\mathsf{T}}^{-}\in\mathbb{R}^{C\times d}$ to refine the negative text embeddings through supervised task-specific training.
-The prediction logit is subsequently given by:
-$\mathsf{T}_{\mathtt{cache}}^{-}\leftarrow\mathsf{Normalize}\left(\mathsf{T}_{\mathtt{cache}}^{-}+\mathcal{R}_{\mathsf{T}}^{-}\right),\quad\mathcal{S}_{\mathsf{T}}^{-}=\delta_{\mathsf{T}}\left(\mathbf{1}-f_{\mathcal{I}}\mathsf{T}_{\mathtt{cache}}^{-\quad\top}\right)\in\mathbb{R}^{1\times C},$ (6)
-where $\delta_{\mathsf{T}}$ is a fixed scaling parameter that adjusts $\mathcal{S}_{\mathsf{T}}^{-}$ to match the mean value of $\mathcal{S}_{\mathsf{T}}^{+}$, and $\mathbf{1}\in\mathbb{R}^{1\times C}$ denotes an all-ones $C$-dimensional vector.
-Negative Visual Adapter.
-In the visual domain, we also similarly pursue some negative image features to enable negative predictions.
-To achieve this, we naturally treat all $(C-1)K$ images from the other $C-1$ classes as negative samples for a given class $c$.
-These negative image samples hold the same property as the negative text descriptions: they are explicitly selected to be non-representative of a specific class, which means that the higher the similarity to them, the lower the probability that the image is classified to that class.
-To mitigate individual biases, we randomly select one image from each of the $C-1$ classes and compute the average of their extracted features to represent the negative image prototypes [56].
-In this way, we can get a total of $K$ negative image prototypes for each of the $C$ classes, thereby constructing a negative visual cache $\mathsf{V}_{\mathtt{cache}}^{-}\in\mathbb{R}^{CK\times d}$.
-Similar to Eq.
-(4), we compute the affinities between the input image and those negative image prototypes, and calculate the classification logit:
-$\mathcal{A}^{-}\!=\!\delta_{\mathsf{V}}\exp\left(-\beta\left(\mathbf{1}-\left(\mathbf{1}-f_{\mathcal{I}}\mathsf{V}_{\mathtt{cache}}^{-\quad\top}\right)\right)\right)$ $\!=\!\delta_{\mathsf{V}}\exp\left(-\beta f_{\mathcal{I}}\mathsf{V}_{\mathtt{cache}}^{-\quad\top}\right)\in\mathbb{R}^{1\times CK},$ (7)
-$\mathcal{S}_{\mathsf{V}}^{-}=\alpha\mathcal{A}^{-}L$ $\in\mathbb{R}^{1\times C}.$ (8)
-where $\delta_{\mathsf{V}}$ is another fixed scaling parameter that adjusts $\mathcal{A}^{-}$/$\mathcal{S}_{\mathsf{V}}^{-}$ to match the mean value of $\mathcal{A}^{+}$/$\mathcal{S}_{\mathsf{V}}^{+}$, $L\in\mathbb{R}^{CK\times C}$ denotes the corresponding one-hot labels.
-To further encode the diverse fine-grained visual knowledge associated with the specific downstream task, we introduce an extra set of learnable residual parameters, $\mathcal{R}_{\mathsf{V}}^{-}\in\mathbb{R}^{C\times d}$, which is similarly broadcast to refine the negative visual cache:
-$\mathsf{V}_{\mathtt{cache}}^{-}\leftarrow\mathsf{Normalize}\left(\mathsf{V}_{\mathtt{cache}}^{-}+\mathcal{R}_{\mathsf{V}}^{-}\right).$ (9)
-3.4 Unified Dual-Path Adapter
-To effectively adapt VLMs to downstream tasks, our DualAdapter combines the functionalities of all four introduced adapters.
-As illustrated in Figure 3 (Left), we provide an overview of the specific example features stored in the cache for the class dog.
-By measuring the similarities between the input image feature and these cached example features, our DualAdapter produces a unified final prediction of the probability that the input image represents a dog.
-The same process is consistently applied across all classes to perform recognition tasks.
-DualAdapter Inference.
-To derive the final classification scores, we aggregate the four outputs from dual-path (i.e., positive and negative) adapters spanning both textual and visual modalities, expressed as:
-$\mathcal{S}_{\mathsf{final}}=\lambda\left(\mathcal{S}_{\mathsf{T}}^{+}+\mathcal{S}_{\mathsf{V}}^{+}\right)+(1-\lambda)\left(\mathcal{S}_{\mathsf{T}}^{-}+\mathcal{S}_{\mathsf{V}}^{-}\right).$ (10)
-Here, $\lambda$ serves as a tuning hyper-parameter to balance the contribution of positive and negative adapter logits.
-Throughout the training process, the collection of learnable parameters =$\mathcal{R}=\{\mathcal{R}_{\mathsf{T}}^{+},\mathcal{R}_{\mathsf{V}}^{+},\mathcal{R}_{\mathsf{T}}^{-},\mathcal{R}_{\mathsf{V}}^{-}\}$ is updated via stochastic gradient descent, employing a cross-entropy loss function for optimization.
-Figure 3: Visualization of cosine similarities on the validation set of ImageNet [4].
-(Left) Example features of the class dog stored in the cache; (Middle) Distribution of pairwise image-text similarities with dual-path features in textual adapters; (Right) Distribution of pairwise image-image similarities with dual-path features in visual adapters.
-Effectiveness of Four Adapters.
-Fig.
-3 qualitatively shows the effectiveness of the four adapters we designed.
-In the middle and right sub-figure, we visualize the distributions of the pairwise cosine similarities between the features of the input image and the textual/visual features stored in the adapters on ImageNet [4] validation set.
-From Fig.
-3, we can observe the following statistical patterns: (1) As we expected, the input image feature is more similar with positive features from the same class (blue $>$ orange) and negative features from other classes (red $>$ green) across both two modalities; (2) Within the textual modality, negative features tend to be less similar to the input image compared to positive features (red $<$ blue, green $<$ orange), since the negative prompts are less common in CLIP [48] training corpus; (3) Within the visual modality, the similarity distribution of negative features occupies an intermediate position between the positive features of the same and different classes (orange $<$ green/red $<$ blue).
-This is because the negative visual features are constructed by averaging the image features across all other classes, inherently leading to a more generic representation that lacks the distinctiveness characteristic of positive features within a single class.
-3.5 Similarity-Based Label Refinement
-Previous adapter-style fine-tuning methods [79, 64] generally assume that the few-shot training samples are meticulously curated to provide accurate representation for each class.
-However, in practical scenarios, this assumption often does not hold true.
-We recognize that not every image is of high quality or equally representative of its respective class, which can be particularly detrimental in the context of few-shot learning, where the availability of training images is extremely limited.
-For example, we provide the t-SNE [38] visualization of few-shot image features from the OxfordPets [45] dataset in Fig.
-4 (Left).
-We argue that the few-shot samples are highly noisy: there is even a data point from the Basset Hound class that overlaps with the prototype (i.e., average feature) of Cocker Spaniel.
-In response to this, we devise an unsupervised similarity-based label refinement process to address this challenge.
-Recall that in our visual adapter, we calculate the image-image affinities between the input image and few-shot training images, with all few-shot images being assigned equal weights [^2] Each few-shot image is assigned a one-hot label, corresponding to a weight of 1.
-as detailed in Eq.
-(4).
-As we discussed, this approach is possibly problematic in the noisy few-shot learning setting.
-Consequently, we extend the visual adapter to assign non-uniform confidence scores according to pairwise similarities, effectively downweighting outliers and emphasizing more representative samples.
-Specifically, our refinement process is based on an intuitive assumption: the representative image feature is closer to other image features from the same class than those low-quality outliers [26].
+Intuitively, if the feature of an input image $f_{v}$ closely resembles the negative textual feature for a specific class $c$, it strongly suggests that the image does not belong to class $c$, i.e., $\mathbb{P}(y=c|\boldsymbol{x})\propto 1-f_{v}^{\top}f_{t_{c}}^{-}$.
+Through supervised task-specific training, we optimize a learnable residue $\mathcal{R}_{\mathsf{T}}^{-}\in\mathbb{R}^{C\times d}$ and obtain the negative predictions based on dissimilarities $1-f_{v}\mathsf{T}_{\mathtt{cache}}^{-}$:
+$\mathsf{T}_{\mathtt{cache}}^{-}\leftarrow\mathsf{Normalize}\left(\mathsf{T}_{\mathtt{cache}}^{-}+\mathcal{R}_{\mathsf{T}}^{-}\right),$
+$\mathcal{S}_{\mathsf{T}}^{-}=\delta_{\mathsf{T}}\left(1-f_{v}\mathsf{T}_{\mathtt{cache}}^{-\quad\top}\right)\in\mathbb{R}^{C},$ (7)
+where $\delta_{\mathsf{T}}$ is a fixed scaling parameter that adjusts $\mathcal{S}_{\mathsf{T}}^{-}$ to match the mean value of $\mathcal{S}_{\mathsf{T}}^{+}$.
+Figure 3: Visualization of cosine similarities on ImageNet [6] validation set.
+We present distributions of pairwise similarities between the input image feature and both the learned positive and negative features from textual (Left) and visual (Right) modalities.
+Negative Visual Branch.
+In the visual domain, we also similarly pursue some negative image features to be non-representative of a specific class, which means that the higher the similarity to them, the lower the probability that the image is classified to that class, i.e., $\mathbb{P}(y=c|\boldsymbol{x})\propto 1-f_{v}^{\top}f_{v_{c}}^{-}$.
+To achieve this, we randomly select one image from each of the $C-1$ classes and compute the average of their extracted features to represent the negative features.
+In this way, we can get a total of $K$ negative visual features for each of the $C$ classes, thereby constructing a negative visual cache $\mathsf{V}_{\mathtt{cache}}^{-}\in\mathbb{R}^{CK\times d}$.
+Symmetric to Eq.
+(6), we compute the negative affinities to obtain the classification logits:
+$\mathsf{V}_{\mathtt{cache}}^{-}\leftarrow\mathsf{Normalize}\left(\mathsf{V}_{\mathtt{cache}}^{-}+\mathcal{R}_{\mathsf{V}}^{-}\right),$
+$\mathcal{S}_{\mathsf{V}}^{-}=\delta_{\mathsf{V}}\mathcal{A}\left(1-f_{v}\mathsf{V}_{\mathtt{cache}}^{-\quad\top}\right)L\in\mathbb{R}^{C}.$ (8)
+where $\delta_{\mathsf{V}}$ is another fixed scaling parameter that adjusts $\mathcal{S}_{\mathsf{V}}^{-}$ to match the mean value of $\mathcal{S}_{\mathsf{V}}^{+}$.
+Here, we also introduce a set of learnable parameters $\mathcal{R}_{\mathsf{V}}^{-}\in\mathbb{R}^{C\times d}$ to refine the negative visual features.
+Final Inference. As discussed in Section 3.1, we ensemble the predictions from both classifiers to derive the final classification scores and predictions:
+$\mathcal{S}_{\mathsf{final}}=\lambda\left(\mathcal{S}_{\mathsf{T}}^{+}+\mathcal{S}_{\mathsf{V}}^{+}\right)+(1-\lambda)\left(\mathcal{S}_{\mathsf{T}}^{-}+\mathcal{S}_{\mathsf{V}}^{-}\right),$
+$\mathbb{P}(y|\boldsymbol{x})=\mathtt{Softmax}\left(\mathcal{S}_{\mathsf{final}}\right).$ (9)
+Here, $\lambda$ serves as a tuning hyper-parameter to balance the contribution of positive and negative logits.
+Throughout the training process, the collection of learnable parameters =$\mathcal{R}=\{\mathcal{R}_{\mathsf{T}}^{+},\mathcal{R}_{\mathsf{V}}^{+},\mathcal{R}_{\mathsf{T}}^{-},\mathcal{R}_{\mathsf{V}}^{-}\}$ is updated via stochastic gradient descent with a cross-entropy loss.
+Qualitative Visualizations.
+Figure 3 qualitatively shows the effectiveness of both the positive and negative classifiers we designed.
+Specifically, we visualize the distributions of pairwise cosine similarities between the input image feature $f_{v}$ and both the learned positive and negative features from two modalities ($f_{t}^{+},f_{t}^{-},f_{v}^{+},f_{v}^{-}$) on the validation set of ImageNet [6].
+We can observe the following statistical patterns: (1) As we expected, the input image feature is more similar with positive features from the same class (blue $>$ orange) and negative features from other classes (red $>$ green) across both two modalities; (2) Within the visual modality, the similarity distribution of negative features occupies an intermediate position between the positive features of the same and different classes (orange $<$ green/red $<$ blue).
+This is because the negative visual features are constructed by averaging the image features across all other classes, inherently leading to a more generic representation that lacks the distinctiveness characteristic of positive features within a single class; (3) Within the textual modality, negative features tend to be less similar to the input image compared to positive features (red $<$ blue, green $<$ orange), since the negative prompts are less common in CLIP [38] training corpus.
+3.5 Few-Shot Instance Reweighting
+In the few-shot adaptation setting, our classifier is trained using only a limited number of samples, with each sample making a significant contribution to the formation of the final decision boundary.
+Consequently, our VLMs are particularly vulnerable to potential noise in the few-shot sample set.
+In Figure 4 (Left), we simulate real-world noise scenarios by randomly flipping the labels of a portion of support samples.
+We demonstrate that the performance of Tip-Adapter-F [71] drastically decreases from 65.52% to 62.47% when the labels of 8 out of the 16 samples per class are randomly flipped.
+Moreover, even if the labels are all correct, we recognize that not every image is of high quality or equally representative of its respective class, as shown in Figure 4 (Right).
+To address this issue, we have developed a few-shot instance reweighting technique to assign non-uniform confidence scores to each sample, effectively downweighting outliers (or mislabelled samples) and prioritizing more representative samples.
+Figure 4: Few-shot instance reweighting.
+(Left) The performance of Tip-Adapter-F [71] degrades drastically when label noise exists in the few-shot sample set; (Right) t-SNE [53] visualization of visual features for 4 random classes from the OxfordPets [35] dataset, where some outliers are marked with red circles.
+Specifically, our proposed instance reweighting is based on an intuitive assumption: the representative image feature is closer to other image features from the same class than those low-quality outliers.
 Based on this assumption, given the $K$-shot image features $\{f_{v_{c}}^{(i)}\}_{i=1}^{K}$ from a specific class $c$, we calculate the average cosine similarities of each image feature to others:
 $\mathsf{d}_{c}^{(i)}\!=\!\frac{1}{K-1}\!\sum_{j,\,j\neq i}\!\cos\left(f_{v_{c}}^{(i)},f_{v_{c}}^{(j)}\right)\!=\!\frac{1}{K-1}\!\sum_{j,\,j\neq i}\!f_{v_{c}}^{(i)\top}\!f_{v_{c}}^{(j)}.$ (10)
-We then assign non-uniform confidence scores $w_{c}^{(i)}$ and compute refined label values $\ell_{c}^{(i)}$ for all $K$-shot image features based on their similarities to others:
+We then assign non-uniform weights $w_{c}^{(i)}$ and compute the reweighted confidences $\ell_{c}^{(i)}$ for all $K$-shot image features based on their average similarities to others:
 $w_{c}^{(i)}=\frac{\exp(\mathsf{d}_{c}^{(i)}/\tau)}{\sum_{i^{\prime}}\exp(\mathsf{d}_{c}^{(i^{\prime})}/\tau)},\quad\ell_{c}^{(i)}=Kw_{c}^{(i)},$ (11)
-where $\tau$ is a temperature hyper-parameter to control the intensity of our refinement.
-In particular, as $\tau\rightarrow\infty$, our refinement becomes inactive, assigning a label value of 1 to all instances.
-Conversely, as $\tau\rightarrow 0$, our refinement process becomes more selective, assigning the highest label value exclusively to the most representative image feature, while all other features are assigned 0.
-This refinement process is applied across each class, wherein the original one-hot labels $L$, are weighted by the new label values computed in Eq.
-(12), resulting in the updated labels $\mathbb{L}$.
-This refinement process is incorporated into both the positive and negative visual adapters; therefore we can rewrite Eqs.
-(4) and (8) as:
-$\mathcal{S}_{\mathsf{V}}^{+}=\alpha\mathcal{A}^{+}\mathbb{L}\in\mathbb{R}^{1\times C},\quad\mathcal{S}_{\mathsf{V}}^{-}=\alpha\mathcal{A}^{-}\mathbb{L}\in\mathbb{R}^{1\times C}.$ (13)
-Figure 4: Motivation for the proposed similarity-based label refinement.
-(Left) t-SNE [38] visualization of visual features for 4 representative classes from the OxfordPets [45] dataset, with prototype [56] of each class denoted by the crossmark $\times$; (Middle) Our refined labels for 4 example images from class Beagle; (Right) An illustrative plot of the modeled distributions using one-hot labels and our refined labels for the Beagle class.
-Effectiveness of Label Refinement.
-In Fig.
-4 (Middle), we present our refined label values for 4 example images from class Beagle in OxfordPets [45] dataset.
-Notably, our unsupervised refinement approach functions by measuring the pairwise similarities among 8-shot images from class Beagle, without requiring any extra parameters or additional training.
-Our similarity-based refinement process allocates higher values to the more representative samples at the top, and lower values to the lower-quality images at the bottom, where the Beagle is either obscured or presented in black and white.
-In Fig.
-4 (Right), we provide an illustration of the modeled visual feature distribution for the Beagle class using one-hot labels and our refined labels.
-Through our refinement process, we are better equipped to capture the underlying feature distribution (typically, a normal distribution [68, 81]) of the Beagle class, whereas the use of one-hot labels restricts us to modeling merely a uniform distribution.
+where $\tau$ is a temperature hyper-parameter to control the intensity of our instance reweighting.
+This process is applied across each class, wherein the original one-hot labels $L$, are reweighted using the new confidence values calculated in Eq.
+(11) to yield $\mathbb{L}$.
+By incorporating this reweighting technique into the visual branches of both classifiers, we can reformulate Eqs.
+(6) and (8) as follows:
+$\mathcal{S}_{\mathsf{V}}^{+}=\mathcal{A}\left(f_{v}\mathsf{V}_{\mathtt{cache}}^{+\quad\top}\right)\mathbb{L}^{+},\,\mathcal{S}_{\mathsf{V}}^{-}=\delta_{\mathsf{V}}\mathcal{A}\left(1-f_{v}\mathsf{V}_{\mathtt{cache}}^{-\quad\top}\right)\mathbb{L}^{-}.$ (12)
 4 Experiments
 In this section, we conduct extensive experiments on two tasks across 15 datasets.
-These results demonstrate that our DualAdapter outperforms other state-of-the-art methods, exhibiting enhanced adaptation and generalization capabilities.
+These results demonstrate that our proposed method is simple yet highly effective, surpassing other state-of-the-art methods in both few-shot adaptation and domain generalization capabilities.
 4.1 Experimental Settings
+Figure 5: Performance comparisons on few-shot learning on 11 image classification datasets.
+For each dataset, we report the mean accuracy and 95% confidence interval over 3 random seeds of our SimNL on 1-/2-/4-/8-/16-shot settings.
 Tasks and Datasets.
-To validate the effectiveness of our proposed DualAdapter, we evaluate our proposed method on two standard benchmarking tasks: few-shot learning and domain generalization, respectively.
+To validate the effectiveness of SimNL, we evaluate our method on two standard benchmarking tasks: few-shot learning and domain generalization, respectively.
 For few-shot learning task, we comprehensively evaluate our method on 11 well-known image classification benchmarks: ImageNet [6], Caltech101 [9], OxfordPets [35], StandfordCars [24], Flowers102 [33], Food101 [2], FGVCAircraft [30], DTD [5], SUN397 [63], EuroSAT [14], and UCF101 [48].
-For domain generalization, we investigate the generalization capability of our DualAdapter on 4 variant datasets of ImageNet: ImageNet-V2 [50], ImageNet-Sketch [65], ImageNet-A [20], and ImageNet-R [19].
+For domain generalization, we evaluate the generalizability of our SimNL on 4 variants of ImageNet: ImageNet-V2 [40], ImageNet-Sketch [55], ImageNet-A [16], and ImageNet-R [15].
+Moreover, we also explore the adaptation of VLMs using a noisy few-shot sample set from ImageNet [6], where labels are randomly flipped to simulate real-world scenarios.
 Implementation Details.
-Following previous works [83, 79], we adopt ResNet-50 [17] backbone as the visual encoder of CLIP in our experiments by default.
-We adopt prompt ensembling, leveraging textual prompts from both CLIP [48] and CuPL [46] to enhance model performance.
-For the negative prompts we used for each dataset, please kindly refer to the Supplementary Materials.
+Following previous works, we adopt ResNet-50 [13] backbone as the visual encoder of CLIP in our experiments by default.
+We adopt prompt ensembling, leveraging textual prompts from CLIP [38] to enhance model performance.
+For the negative prompts we used for each dataset, please kindly refer to Appendix C.2.
 We set the hyper-parameters $\lambda$ and $\tau$ as $0.75$ and $1$, respectively.
-Our DualAdapter is trained using the AdamW [36] optimizer with a cosine scheduler [35].
+Our SimNL is trained using the AdamW [29] optimizer with a cosine scheduler [28].
 The batch size is set to 256.
 For $\mathcal{R}_{\mathsf{T}}^{+}$ and $\mathcal{R}_{\mathsf{V}}^{+}$, the learning rate is set to $0.0001$, while for $\mathcal{R}_{\mathsf{T}}^{-}$ and $\mathcal{R}_{\mathsf{V}}^{-}$, the learning rate is set to $0.0005$.
-Additionally, our model is trained for 200 epochs on the EuroSAT [18] dataset, and for 20 epochs on all other datasets.
+Our model is trained for 200 epochs on the EuroSAT [14] dataset, and for 20 epochs on all other datasets.
 To ensure the reliability of our results, we perform each experiment three times using different initialization seeds and report the mean accuracy achieved.
-All experiments are conducted on a single NVIDIA RTX 6000 Ada GPU.
+All experiments are conducted on a single 48GB NVIDIA RTX 6000 Ada GPU.
 Baselines.
@@
 For a fair comparison, we directly report the results of these baselines from their respective original papers.
-4.2 Few-Shot Learning
-In Fig.
-5, we compare the few-shot learning performance of our DualAdapter with other state-of-the-art methods on 11 image classification datasets.
+4.2 Results and Analysis
+Few-Shot Learning.
+In Figure 5, we compare the few-shot learning performance of our proposed method with other state-of-the-art methods on 11 image classification datasets.
 In the top-left sub-figure, we also present the average classification accuracy across all 11 datasets.
-The results indicate that our DualAdapter consistently outperforms other methods across various few-shot learning settings by substantial margins.
-Moreover, our DualAdapter demonstrates more pronounced performance improvements in specialized classification tasks, such as satellite image and texture classification on the EuroSAT [18] and DTD [3] datasets.
-With 16-shot training on these two datasets, our DualAdapter surpasses Tip-Adapter-F [79] by a notable 3.56% and 3.50%, and enhances the performance of zero-shot CLIP [48] significantly by 45.56% and 30.50%, respectively.
-For full numerical results, please refer to our Supplementary Materials.
-Overall, the consistently superior performance on 11 datasets fully demonstrates the effectiveness of our DualAdapter.
-Figure 5: Performance comparisons on few-shot learning on 11 image classification datasets.
-For each dataset, we report the mean accuracy and 95% confidence interval over 3 random seeds of our DualAdapter on 1-/2-/4-/8-/16-shot settings.
-4.3 Robustness to Natural Distribution Shifts
+The results indicate that our method consistently outperforms other methods across various few-shot learning settings by substantial margins.
+Moreover, our proposed method demonstrates more pronounced performance improvements in specialized classification tasks, such as satellite image and texture classification on the EuroSAT [14] and DTD [5] datasets.
+With 16-shot training on these two datasets, our method surpasses Tip-Adapter-F [71] by a notable 3.56% and 3.50%, respectively.
+For full numerical results, please refer to Table A1 in Appendix A.1.
+Overall, the consistently superior performance on 11 datasets fully demonstrates the general effectiveness of our proposed approach.
+Robustness to Natural Distribution Shifts.
+In Table 1, we compare the generalizability of our SimNL with other methods in the presence of distribution shifts.
+Specifically, all the models are trained solely on 16-shot ImageNet [6], and directly tested on 4 out-of-distribution ImageNet variant datasets.
+As shown in Table 1, our method not only achieves state-of-the-art performance on the source dataset but also attains an average performance gain of 1.18% across 4 out-of-distribution (OOD) target datasets.
+These experimental results indicate that by enabling adaptation from both positive and negative perspectives, our method enhances robustness against distribution shifts.
 Table 1: Performance comparison on robustness to distribution shifts.
@@
 GraphAdapter [25] 64.94 56.58 35.89 23.07 60.86 44.10
-DualAdapter (Ours) 66.52 57.87 36.38 25.73 61.12 45.28
-In Table 1, we compare the generalization performance of our DualAdapter with other state-of-the-art methods in the presence of distribution shifts.
-Specifically, all the models are trained solely on 16-shot ImageNet [4], and directly tested on 4 out-of-distribution ImageNet variant datasets.
-As shown in Table 1, our DualAdapter not only achieves state-of-the-art performance on the source dataset but also attains an average performance gain of 1.18% across 4 out-of-distribution (OOD) target datasets.
-These experimental results demonstrate that by enabling adaptation through both positive and negative paths, our DualAdapter can more effectively prevent overfitting to the source dataset, thereby enhancing robustness against distribution shifts.
-4.4 Ablation Studies
-[…]
-Table 2: Ablation studies for different variants of DualAdapter. We evaluate the few-shot adaptation capabilities of four DualAdapter variants on ImageNet [4].
-# Method $\mathcal{R}_{\mathsf{T}}^{+}$ $\mathcal{R}_{\mathsf{T}}^{-}$ $\mathcal{R}_{\mathsf{V}}^{+}$ $\mathcal{R}_{\mathsf{V}}^{-}$ 1-shot 2-shot 4-shot 8-shot 16-shot
-[…]
-[…]
-3 ${}^{+}$ 62.83 63.31 63.95 65.13 66.27
-4 ${}^{-}$ 62.65 63.07 63.60 64.36 65.12
-5 DualAdapter 62.89 63.47 64.12 65.37 66.52
-Effects of Different Visual Backbones.
-We also implement our DualAdapter with various visual encoders, including ResNet [17] and ViT [6], and evaluate its performance against other fine-tuning approaches in Fig.
-6 (Left).
-We can see that our DualAdapter consistently exceeds other methods across all visual backbones, indicating the general effectiveness of our approach.
+SimNL (Ours) 66.52 57.87 36.38 25.73 61.12 45.28
+Robustness to Label Noise.
+In Table 2, we report the performance of Tip-Adapter-F [71] and our SimNL on a noisy 16-shot ImageNet [6], where we randomly flip 10% to 50% labels.
+As shown, applying our instance reweighting enhances both Tip-Adapter-F and SimNL’s robustness to label noise.
+Notably, our reweighting technique also improves performance when no label noise is introduced, as it can identify and downweight outliers during adaptation.
+Table 2: Comparison of robustness to label noise on noisy 16-shot ImageNet [6].
+We apply our instance reweighting technique to both Tip-Adapter-F [71] and our SimNL, and report the performance across four levels of noise.
+Method 0% 10% 25% 50%
+Tip-Adapter-F [71] 65.52 64.93 64.04 62.47
++ Reweighting 65.64 65.25 64.55 63.39
+Performance Gain +0.12 +0.32 +0.51 +0.92
+SimNL (Ours) 66.31 65.54 64.77 63.37
++ Reweighting 66.52 65.82 65.16 64.02
+Performance Gain +0.21 +0.28 +0.39 +0.65
 Figure 6: More ablation results.
-(Left) Performance comparison of our DualAdapter with other methods on few-shot learning task using different visual backbones; (Middle) Sensitivity analysis of $\lambda$ from Eq.
+(Left) Performance comparison of our SimNL with others on few-shot learning using different visual backbones; (Middle) Sensitivity analysis of $\lambda$ from Eq.
 (9) on ImageNet [6]; (Right) Sensitivity analysis of $\tau$ from Eq.
-(12) on 3 datasets, with the value of $\tau$ presented on a logarithmic scale.
-[…]
-Scalability to More Shots.
-When scaling up to more than 16 shots, Tip-Adapter-F [79] limits the cache size to prevent memory overflow.
-In contrast, our DualAdapter features reduced memory requirements, enabling it to operate without imposing cache size constraints, even in the 128-shot setting.
-Performance comparisons in Table 4.4 indicate that our DualAdapter outperforms Tip-Adapter-F [79] more significantly in higher shot settings, e.g., by 1.24% in the 128-shot setting.
+(11) on 3 datasets.
 Efficiency Comparison.
-We also compare the efficiency of DualAdapter with existing methods in Table 4.4.
-Our DualAdapter achieves the highest accuracy while also exhibiting advantageous computational efficiency: (1) Compared to prompt-based learning methods such as CoOp [84] and ProGrad [87], our DualAdapter requires approximately 300$\times$ less training time and over 1000$\times$ fewer FLOPs since we do not need to propagate gradients through the textual encoder; (2) Compared to adapter-style fine-tuning methods, our DualAdapter is also competitive in efficiency.
-It requires 10$\times$ less training time than CLIP-Adapter [12], and demands 3$\times$ fewer FLOPs and 4$\times$ fewer learnable parameters than Tip-Adapter-F [79].
-Table 3: Performance comparison with Tip-Adapter(-F) [79] on ImageNet [4] with 16-/32-/64-/128-shot settings.
-Table 4: Efficiency comparison with other existing methods on 16-shot ImageNet [4].
-Method 16-shot 32-shot 64-shot 128-shot
-CLIP [48] ${}^{\dagger}$0-shot: 60.33
-Tip-Adapter [79] 62.03 62.51 62.88 63.15
-Tip-Adapter-F [79] 65.47 66.58 67.96 69.74
-DualAdapter (Ours) 66.52 67.68 69.01 70.98
+We also compare the efficiency of SimNL with existing methods in Table 3.
+Our method achieves the highest accuracy while also exhibiting advantageous computational efficiency: (1) Compared to prompt-based learning methods such as CoOp [74] and ProGrad [75], our proposed method requires approximately 300$\times$ less training time and over 1000$\times$ fewer FLOPs since we do not need to propagate gradients through the textual encoder; (2) Compared to adapter-style fine-tuning methods, our SimNL requires 10$\times$ less training time than CLIP-Adapter [10], and demands 3$\times$ fewer FLOPs than Tip-Adapter-F [71] and GraphAdapter [25].
+Table 3: Efficiency comparison with other existing methods on 16-shot ImageNet [6]. We report the training time, number of epochs, FLOPs, and the number of parameters for each method.
 Method Training Epochs GFLOPs Param. Accuracy
@@
 Tip-Adapter-F [71] 5 min 20 0.030 16.38M 65.51
-DualAdapter (Ours) 5 min 20 0.009 4.10M 66.52
-Table 4: Efficiency comparison with other existing methods on 16-shot ImageNet [4].
+GraphAdapter [25] 5 min 20 0.030 4.15M 65.70
+SimNL (Ours) 5 min 20 0.009 4.15M 66.52
+4.3 Ablation Studies
+Effectiveness of Different Components.
+In Table 4, we conduct a systematic analysis of the impacts of various components within our SimNL framework.
+More specifically, we assess the performance of four distinct SimNL variants, each configured to allow two learnable residues to be updated while keeping the others fixed.
+We observe that the textual variant (SimNL-T) and the positive variant (SimNL-P) generally exhibit superior efficiency compared to their visual and negative counterparts.
+The full SimNL method, which integrates all four branches, surpasses the individual variants by achieving the highest average accuracy of 77.80% in the 16-shot scenario.
+Table 4: Ablation studies for different variants of our method. We evaluate the few-shot adaptation capabilities of four variants and report their average performance across all 11 datasets.
+Method $\mathcal{R}_{\mathsf{T}}^{+}$ $\mathcal{R}_{\mathsf{T}}^{-}$ $\mathcal{R}_{\mathsf{V}}^{+}$ $\mathcal{R}_{\mathsf{V}}^{-}$ 1-shot 2-shot 4-shot 8-shot 16-shot
+SimNL-T 67.08 69.70 71.86 74.17 77.08
+SimNL-V 64.60 66.03 68.25 71.84 74.41
+SimNL-P 66.72 69.30 71.49 73.98 76.76
+SimNL-N 66.23 68.10 70.55 72.33 74.17
+SimNL 67.45 70.06 72.43 75.06 77.80
+Effects of Different Visual Backbones.
+We also implement our SimNL with various visual encoders, including ResNet [13] and ViT [8], and evaluate their performance against other adaptation methods in Figure 6 (Left).
+We can see that our SimNL consistently exceeds other methods across all visual backbones, indicating the general effectiveness of our negative learning approach.
+Sensitivity Analysis of Hyper-Parameters.
+We provide sensitivity analysis for the hyper-parameters $\lambda$ and $\tau$ in Figure 6.
+The hyper-parameter $\lambda$ from Eq.
+(9) controls the combination of the positive and negative predictions.
+In Figure 6 (Middle), we can observe that setting $\lambda=0.75$ consistently yields the optimal performance.
+Moreover, in Figure 6 (Right), we adjust the temperature hyper-parameter $\tau$ in Eq.
+(11) from $0.04$ to $25$ and report the 16-shot accuracy of our method.
+We can see that by reweighting the few-shot samples, we achieve performance improvements ranging from 0.21% to 0.55% across three datasets.
+Scalability to More Shots.
+In Table 5, we demonstrate that our SimNL scales effectively to more-shot settings, consistently outperforming Tip-Adapter-F [71] by 1.05% to 1.24% as the number of shots increases.
+Table 5: Scalability to more shots. We show performance comparison with Tip-Adapter(-F) on ImageNet at higher-shot settings.
+Method 16-shot 32-shot 64-shot 128-shot
+CLIP [38] 0-shot: 60.33
+Tip-Adapter [71] 62.03 62.51 62.88 63.15
+Tip-Adapter-F [71] 65.47 66.58 67.96 69.74
+SimNL (Ours) 66.52 67.68 69.01 70.98
 5 Conclusion
-In this work, we introduce DualAdapter, a novel approach to effectively adapt vision-language models for downstream tasks by integrating both positive and negative adapters.
-This innovative approach allows for effective few-shot adaptation from a dual perspective.
-Furthermore, we develop an unsupervised similarity-based label refinement approach to mitigate the adverse effects of low-quality image samples during the few-shot adaptation process.
-Comprehensive evaluations on 15 diverse datasets demonstrate that DualAdapter outperforms the state-of-the-art methods in both few-shot learning and domain generalization tasks.
-Limitations.
-We identify two potential limitations of our DualAdapter: (1) Its efficacy in zero-shot situations is constrained due to the scarcity of original negative descriptions in the CLIP training corpus; (2) Similar to other adapter-style approaches such as Tip-Adapter-F [79], our DualAdapter fine-tuned on a specific task cannot be directly applied to another task without additional adaptation.
-However, recently, Wang et al.
-[68] shows that adapter-style fine-tuning methods can be extended for these scenarios using the $k$NN algorithm.
-Acknowledgement
-This work has been funded in part by the Army Research Laboratory (ARL) under grant W911NF-23-2-0007 and W911NF-19-2-0146, and the Air Force Office of Scientific Research (AFOSR) under grants FA9550-18-1-0097 and FA9550-18-1-0251.
+In this work, we introduce SimNL, a simple and effective approach that applies the concept of negative learning to vision-language few-shot adaptation.
+Specifically, we transform the open-vocabulary CLIP model into a negative classifier, which is further optimized to exclude incorrect classes based on the image input.
+During inference, we conduct simultaneous positive classification and negative exclusion to enhance the overall prediction accuracy.
+Furthermore, we develop an unsupervised few-shot instance reweighting approach to mitigate the adverse effects of noisy image samples during few-shot adaptation .
+Comprehensive evaluations on 15 diverse datasets demonstrate that our proposed SimNL outperforms the state-of-the-art methods in both few-shot learning and domain generalization tasks, while maintaining competitive efficiency.
+Acknowledgements. This work has been funded in part by the Army Research Laboratory (ARL) award W911NF-23-2-0007, DARPA award FA8750-23-2-1015, and ONR award N00014-23-1-2840.
 Supplementary Material
-In this supplementary document, we provide additional details and experimental results to enhance understanding and insights into our proposed DualAdapter.
-This supplementary document is organized as follows:
+In appendix, we provide additional details and experimental results to enhance understanding and insights into our proposed SimNL. This supplementary document is organized as follows:
   • $\bullet$
@@
   • $\bullet$
-    We compare our DualAdapter with other state-of-the-art methods in domain generalization tasks, utilizing an alternative ViT backbone, in Section A.2.
+    We compare our SimNL with other state-of-the-art methods in domain generalization tasks, utilizing an alternative ViT backbone, in Section A.2.
   • $\bullet$
@@
   • $\bullet$
-    Detailed statistics for all utilized datasets are provided in Section B.1.
+    We differentiate our method from related work in Section B.
   • $\bullet$
-    We list the specific positive and negative prompts we used for each dataset in Section B.2.
+    Detailed statistics for all utilized datasets are provided in Section C.1.
   • $\bullet$
-    Finally, we explore potential avenues for future research in Section C.
+    We present the specific positive and negative prompts we used for each dataset in Section C.2.
+  • $\bullet$
+    We list the license information for all used assets in Section D.
+  • $\bullet$
+    Finally, we explore potential future work and discuss the limitations and broader impacts of this work in Section E.
 A Additional Experimental Results
 A.1 Full Numerical Results on Few-Shot Learning
-In Fig.
-5 in Section 4.2 of the main text, we have evaluated our DualAdapter on few-shot learning task and compared with other state-of-the-art methods.
+In Figure 5 of the main text, we have evaluated our SimNL on few-shot learning task and compared with other state-of-the-art methods.
 In Table A1, we present the corresponding full numerical results on the few-shot learning task.
-We also report the 95% confidence interval over 3 random seeds of our DualAdapter to ensure reliability of our results.
+We also report the 95% confidence interval over 3 random seeds of our SimNL to ensure reliability of our results.
 In the last column, we present the average recognition accuracy over 11 datasets.
-The results indicate that our DualAdapter consistently outperforms other state-of-the-art methods across various few-shot learning settings by substantial margins.
-Our DualAdapter demonstrates superior recognition performance across nearly all tested scenarios, with certain exceptions in lower-shot settings for the Food101 [1] and OxfordPets [45] datasets.
+The results indicate that our SimNL consistently outperforms other state-of-the-art methods across various few-shot learning settings by substantial margins.
+Our SimNL demonstrates superior recognition performance across nearly all tested scenarios, with certain exceptions in lower-shot settings for the Food101 [2] and OxfordPets [35] datasets.
 We attribute this to the prevalent challenge of overfitting, a common issue not exclusive to our approach but also affecting many existing methods, especially TaskRes [65]: While TaskRes consistently secures the second-best performance across the other 9 datasets, it underperforms significantly on these two.
@@
 However, in this work, we have designed a label refinement mechanism specifically to address this issue.
-As a result, our DualAdapter secures a relatively robust performance and achieves the second-best on these two datasets.
+As a result, our SimNL secures a relatively robust performance and achieves the second-best on these two datasets.
 Table A1: Full numerical results on few-shot learning task.
-For each dataset, we report the mean accuracy and 95% confidence interval over 3 random seeds of our DualAdapter on 1-/2-/4-/8-/16-shot settings.
-${}^{\dagger}$e report the zero-shot performance of CLIP [48] for all settings.
+For each dataset, we report the mean accuracy and 95% confidence interval over 3 random seeds of our SimNL on 1-/2-/4-/8-/16-shot settings.
+We report the zero-shot performance of CLIP [38] for all settings.
 For TaskRes [65], we report the results using the enhanced base classifier (i.e., TaskRes*).
@@
 Avg.
-${}^{\dagger}$ero-shot CLIP [48] 1-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
+Zero-shot CLIP [38] 1-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
 CoOp [74] 87.43 44.13 50.51 9.80 67.90 73.71 57.15 86.51 55.48 60.10 62.10 59.53
@@
 90.87 53.13 67.70 23.61 84.17 77.93 62.89 86.90 61.34 65.13 68.25 67.45
-DualAdapter (Ours) (±$\pm 0.33$) (±$\pm 0.48$) (±$\pm 0.47$) (±$\pm 0.18$) (±$\pm 0.34$) (±$\pm 0.22$) (±$\pm 0.13$) (±$\pm 0.31$) (±$\pm 0.28$) (±$\pm 0.08$) (±$\pm 0.23$) (±$\pm 0.28$)
-${}^{\dagger}$ero-shot CLIP [48] 2-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
+SimNL (Ours) (±$\pm 0.33$) (±$\pm 0.48$) (±$\pm 0.47$) (±$\pm 0.18$) (±$\pm 0.34$) (±$\pm 0.22$) (±$\pm 0.13$) (±$\pm 0.31$) (±$\pm 0.28$) (±$\pm 0.08$) (±$\pm 0.23$) (±$\pm 0.28$)
+Zero-shot CLIP [38] 2-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
 CoOp [74] 87.92 45.04 60.43 18.25 77.47 72.26 55.88 82.36 58.10 59.82 64.13 61.97
@@
 91.19 59.17 74.40 26.22 88.43 78.35 63.47 87.68 64.77 66.73 70.25 70.06
-DualAdapter (Ours) (±$\pm 0.26$) (±$\pm 0.31$) (±$\pm 0.87$) (±$\pm 0.37$) (±$\pm 0.18$) (±$\pm 0.17$) (±$\pm 0.06$) (±$\pm 0.28$) (±$\pm 0.21$) (±$\pm 0.33$) (±$\pm 0.48$) (±$\pm 0.31$)
-${}^{\dagger}$ero-shot CLIP [48] 4-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
+SimNL (Ours) (±$\pm 0.26$) (±$\pm 0.31$) (±$\pm 0.87$) (±$\pm 0.37$) (±$\pm 0.18$) (±$\pm 0.17$) (±$\pm 0.06$) (±$\pm 0.28$) (±$\pm 0.21$) (±$\pm 0.33$) (±$\pm 0.48$) (±$\pm 0.31$)
+Zero-shot CLIP [38] 4-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
 CoOp [74] 89.17 53.38 70.20 21.72 85.81 72.72 59.93 87.22 61.92 63.46 67.08 66.60
@@
 92.21 66.01 76.54 28.95 92.04 78.74 64.12 88.13 67.96 68.59 73.46 72.43
-DualAdapter (Ours) (±$\pm 0.21$) (±$\pm 0.49$) (±$\pm 0.59$) (±$\pm 0.29$) (±$\pm 0.25$) (±$\pm 0.11$) (±$\pm 0.17$) (±$\pm 0.26$) (±$\pm 0.36$) (±$\pm 0.29$) (±$\pm 0.51$) (±$\pm 0.32$)
-${}^{\dagger}$ero-shot CLIP [48] 8-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
+SimNL (Ours) (±$\pm 0.21$) (±$\pm 0.49$) (±$\pm 0.59$) (±$\pm 0.29$) (±$\pm 0.25$) (±$\pm 0.11$) (±$\pm 0.17$) (±$\pm 0.26$) (±$\pm 0.36$) (±$\pm 0.29$) (±$\pm 0.51$) (±$\pm 0.32$)
+Zero-shot CLIP [38] 8-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
 CoOp [74] 90.15 59.88 76.51 25.93 90.84 71.52 60.91 86.40 68.49 65.63 71.81 69.82
@@
 93.40 67.78 81.62 33.90 95.23 79.23 65.37 89.29 72.08 70.93 76.84 75.06
-DualAdapter (Ours) (±$\pm 0.23$) (±$\pm 0.55$) (±$\pm 0.54$) (±$\pm 0.63$) (±$\pm 0.26$) (±$\pm 0.21$) (±$\pm 0.09$) (±$\pm 0.21$) (±$\pm 0.51$) (±$\pm 0.15$) (±$\pm 0.39$) (±$\pm 0.34$)
-${}^{\dagger}$ero-shot CLIP [48] 16-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
+SimNL (Ours) (±$\pm 0.23$) (±$\pm 0.55$) (±$\pm 0.54$) (±$\pm 0.63$) (±$\pm 0.26$) (±$\pm 0.21$) (±$\pm 0.09$) (±$\pm 0.21$) (±$\pm 0.51$) (±$\pm 0.15$) (±$\pm 0.39$) (±$\pm 0.34$)
+Zero-shot CLIP [38] 16-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
 CoOp [74] 91.61 63.11 82.36 31.01 94.39 73.80 62.95 87.30 72.51 69.11 75.70 73.07
@@
 93.77 70.83 87.36 40.27 96.51 79.87 66.52 90.58 77.48 72.32 80.28 77.80
-DualAdapter (Ours) (±$\pm 0.33$) (±$\pm 0.77$) (±$\pm 0.84$) (±$\pm 0.53$) (±$\pm 0.58$) (±$\pm 0.34$) (±$\pm 0.13$) (±$\pm 0.38$) (±$\pm 0.68$) (±$\pm 0.24$) (±$\pm 0.26$) (±$\pm 0.50$)
+SimNL (Ours) (±$\pm 0.33$) (±$\pm 0.77$) (±$\pm 0.84$) (±$\pm 0.53$) (±$\pm 0.58$) (±$\pm 0.34$) (±$\pm 0.13$) (±$\pm 0.38$) (±$\pm 0.68$) (±$\pm 0.24$) (±$\pm 0.26$) (±$\pm 0.50$)
 A.2 More Results on Domain Generalization
@@
 GraphAdapter [25] 68.47 59.10 42.70 31.73 69.43 50.74
-DualAdapter (Ours) 69.63 59.76 43.41 32.48 69.60 51.31
-In Table 1 in Section 4.3 of the main text, we compare the generalization performance of our DualAdapter with other state-of-the-art methods in the presence of distribution shifts with ResNet-50 visual backbone.
+SimNL (Ours) 69.63 59.76 43.41 32.48 69.60 51.31
+In Table 1 in Section 4.2 of the main text, we compare the generalization performance of our SimNL with other state-of-the-art methods in the presence of distribution shifts with ResNet-50 visual backbone.
 In Table A2, we present the performance comparison on domain generalization task using ViT-B/32 visual backbone.
-Consistently, our DualAdapter not only achieves state-of-the-art performance on the source dataset but also attains an average performance gain of 0.57% across 4 out-of-distribution (OOD) target datasets.
-This verifies that our DualAdapter demonstrates superior generalizability compared to other state-of-the-art methods, independent of the visual backbone utilized.
+Consistently, our SimNL not only achieves state-of-the-art performance on the source dataset but also attains an average performance gain of 0.57% across 4 out-of-distribution (OOD) target datasets.
+This verifies that our SimNL demonstrates superior generalizability compared to other state-of-the-art methods, independent of the visual backbone utilized.
 A.3 More Sensitivity Analyses of Hyper-Parameters
 Table A3: Sensitivity of hyper-parameters. All the results are reported on 16-shot ImageNet [6].
-$\alpha$ 0.0 0.5 1.0 1.2 1.5 2.0 66.34 66.40 66.47 66.52 66.44 66.36 $\beta$ 1.0 1.5 2.0 2.5 3.0 3.5 66.38 66.44 66.52 66.50 66.48 66.40
+$\alpha$ 0.0 0.5 1.0 1.2 1.5 2.0 66.14 66.30 66.41 66.52 66.44 66.36 $\beta$ 1.0 1.5 2.0 2.5 3.0 3.5 66.38 66.44 66.52 66.50 66.48 66.40
 Building upon the sensitivity analyses of $\lambda$ and $\tau$ detailed in Section 4.3 of the main text, this section extends our examination to include the sensitivity of parameters $\alpha$ and $\beta$ on 16-shot ImageNet [6].
@@
 We can see that our choice of $\alpha=1.2$ and $\beta=2.0$ yields the highest performance.
-Moreover, our DualAdapter maintains robust performance when adjusting these two hyper-parameters, since our DualAdapter includes adapters from both textual and visual modalities and each of our four adapters can work effectively, as we presented in Table 2 in the main text.
-B Additional Implementation Details
-B.1 Dataset Details
-In Table B4, we present the detailed statistics of each dataset we used in our experiments, including the number of classes, the sizes of training, validation and testing sets, and their original tasks.
-Table B4: Detailed statistics of datasets used in experiments. Note that the last 4 ImageNet variant datasets are designed for evaluation and only contain the test sets.
+Moreover, our SimNL maintains robust performance when adjusting these two hyper-parameters, since our SimNL includes adapters from both textual and visual modalities and each of our four adapters can work effectively, as we presented in Table 4 in the main text.
+B More Discussions on Related Work
+B.1 Differences between Our Proposed Method and Traditional Negative Learning
+In this work, we apply the concept of negative learning to vision-language few-shot adaptation.
+However, our method is different with traditional negative learning approaches [20, 66].
+Specifically, traditional negative learning approaches aim to learn a positive classifier using negatibe labels, while our method aim to learn a negative classifier using the positive label.
+Negative Learning.
+In the negative learning setting [20, 66], we are given complementary labels $\bar{y}\in\mathcal{Y}\,\backslash\{y\}$, which represents the image does not belong to a specific class.
+Similarly, the encoded one-hot labels are given by $\bar{\boldsymbol{y}}\in\{0,1\}^{C}$.
+This leads to a complementary cross-entropy loss for optimizing the classifier parameters $\theta$.
+The training objective for $\mathcal{F}_{\theta}$ can thus be expressed as
+$\min_{\theta}\mathcal{R}(\mathcal{F}_{\theta})=\mathbb{E}_{(\boldsymbol{x},\bar{y})\sim P(\boldsymbol{x},\bar{y})}\left[\mathcal{L}\left(\mathcal{F}_{\theta}(\boldsymbol{x}),\bar{y}\right)\right],$
+$\text{where}\,\,\mathcal{L}\left(\mathcal{F}_{\theta}(\boldsymbol{x}),\bar{y}\right)=-\sum_{k=1}^{C}\bar{\boldsymbol{y}}_{k}\log(1-\boldsymbol{p}_{k}).$ (13)
+Here, the goal is to minimize the probability corresponding to the complementary label, i.e., $\boldsymbol{p}_{\bar{y}}\rightarrow 0$.
+Our Method.
+In this work, we apply the concept of negative learning to vision-language few-shot adaptation by employing a distinct CLIP-based negative classifier $\mathcal{G}_{\varphi}:\mathcal{X}\rightarrow\mathbb{R}^{C}$ with parameters $\varphi$.
+This classifier predicts the negative probability $\bar{\boldsymbol{p}}=\texttt{Softmax}(\mathcal{G}_{\varphi}(\boldsymbol{x}))$ that the image does not belong to specific classes.
+The expected classification risk for $\mathcal{G}_{\varphi}$ can be written as
+$\min_{\varphi}\mathcal{R}(\mathcal{G}_{\varphi})=\mathbb{E}_{(\boldsymbol{x},y)\sim P(\boldsymbol{x},y)}\left[\mathcal{L}\left(\mathcal{G}_{\varphi}(\boldsymbol{x}),y\right)\right],$
+$\text{where}\,\,\mathcal{L}\left(\mathcal{G}_{\varphi}(\boldsymbol{x}),y\right)=-\sum_{k=1}^{C}\boldsymbol{y}_{k}\log(1-\bar{\boldsymbol{p}}_{k}).$ (14)
+By optimizing this risk, we aim to reduce the negative probability $\bar{\boldsymbol{p}}_{y}\rightarrow 0$ for the true label.
+B.2 Differences between Our Proposed Method and Contrastive Learning
+In this work, we propose a negative learning-based approach for vision-language few-shot adaptation.
+However, this “negative" refers to a negative classifier, not the negative samples pairs in contrastive learning literature.
+In the following discussions, we show that our negative learning is fundamentally different with contrastive learning.
+Specifically, CLIP operates as a similarity-based classifier where classification logits are derived from the similarities between image features $f_{v}$ and class-specific text features $\{f_{t_{c}}\}_{c=1}^{C}$:
+$\boldsymbol{p}_{y}=\frac{\exp\left(\mathrm{cos}\left(f_{t_{y}},f_{v}\right)/t\right)}{\sum\nolimits_{t^{\prime}}{\exp\left(\mathrm{cos}\left(f_{t^{\prime}},f_{v}\right)/t\right)}},$ (15)
+In the context of few-shot adaptation of CLIP, both prompt-based learning methods and adapter-style fine-tuning methods seek to optimize $\{f_{t_{c}}\}_{c=1}^{C}$, whether by tuning the input prompt or directly modulating these features.
+As discussed in Section 3.1, these methods primarily employ positive learning, updating parameters via cross-entropy loss:
+$\mathcal{L}\left(\mathcal{F}_{\theta}(\boldsymbol{x}),y\right)$ $=-\sum_{k=1}^{C}\boldsymbol{y}_{k}\log\boldsymbol{p}_{k}$
+$=-\log\frac{\exp\left(\mathrm{cos}\left(f_{t_{y}},f_{v}\right)/t\right)}{\sum\nolimits_{t^{\prime}}{\exp\left(\mathrm{cos}\left(f_{t^{\prime}},f_{v}\right)/t\right)}},$ (16)
+where $y$ is the ground-truth label for sample $\boldsymbol{x}$.
+This update strategy aims to optimize $p_{y}=1$.
+By considering $f_{t_{y}}$ and $f_{v}$ as a positive pair and $f_{t_{\neq y}}$ and $f_{v}$ as negative pairs, the positive learning process essentially becomes contrastive in nature.
+Specifically, Eq.
+(B.2) has the same form with the contrastive InfoNCE loss [34], which seeks to minimize the distance between the positive pair $f_{t_{y}}$ and $f_{v}$ while maximizing the distance between the negative pairs $f_{t_{\neq y}}$ and $f_{v}$.
+However, in this work, we apply negative learning and introduce another negative classifier $\mathcal{G}_{\varphi}$.
+Specifically, we mine a general negative feature $f_{t_{c}}^{-}$ for each class $c$, which is absent in samples from class $c$ but present in samples from all other classes.
+Then the probability that the image not belongs to class $y$ can be written as
+$\bar{\boldsymbol{p}}_{y}=\frac{\exp\left(\mathrm{cos}\left(f_{t_{y}}^{-},f_{v}\right)/t\right)}{\sum\nolimits_{t^{\prime}}{\exp\left(\mathrm{cos}\left(f_{t^{\prime}}^{-},f_{v}\right)/t\right)}}$ (17)
+As discussed in Section 3.1, we update the set of negative features by
+$\mathcal{L}\left(\mathcal{G}_{\varphi}(\boldsymbol{x}),y\right)$ $=-\sum_{k=1}^{C}\boldsymbol{y}_{k}\log(1-\bar{\boldsymbol{p}}_{k})$
+$=-\log\left(1-\frac{\exp\left(\mathrm{cos}\left(f_{t_{y}}^{-},f_{v}\right)/t\right)}{\sum\nolimits_{t^{\prime}}{\exp\left(\mathrm{cos}\left(f_{t^{\prime}}^{-},f_{v}\right)/t\right)}}\right).$ (18)
+In the language of contrastive learning, we consider all $f_{t_{\neq y}}^{-}$ and $f_{v}$ as positive pairs and $f_{t_{y}}^{-}$ and $f_{v}$ as the only negative pair.
+In summary, the concept of “negative" is different in our negative learning and contrastive learning: (1) In our negative learning, “negative" specifically refers to a negative classifier.
+We explicitly train another negative classifier $\mathcal{G}_{\varphi}$ to ensemble with the positive classifier; (2) In contrastive learning, “negative" refers to the negative sample pairs that are constructed and utilized during training.
+Specifically, by constructing positive and negative sample pairs, the objective of each classifier can be formulated as a contrastive objective as discussed above.
+B.3 Differences with Other Similar Approaches
+To the best of our knowledge, the following works adopt similar negative learning methods to open-set problems:
+  • RPL [3] is the seminal work that learns a set of reciprocal points as negative representations of target classes to enhance model’s recognition capabilities when handling unseen samples.
+  • CLIPN [56] leverages additional large-scale datasets to a fine-tune a “no" text encoder to enhance CLIP’s out-of-distribution (OOD) detection capability.
+  • LSN [32] also focus on learning some negative prompts for OOD detection task using CLIP.
+However, we formulate negative learning from a different and more fundamental perspective, i.e., using negative prediction to directly improve the accuracy of positive prediction in a closed-set problem.
+We also empirically validate that leveraging negative cues from CLIP can effectively improve both few-shot classification performance and generalization capability.
+C Additional Technical Details
+C.1 Dataset Details
+In Table C4, we present the detailed statistics of each dataset we used in our experiments, including the number of classes, the sizes of training, validation and testing sets, and their original tasks.
+Table C4: Detailed statistics of datasets used in experiments. Note that the last 4 ImageNet variant datasets are designed for evaluation and only contain the test sets.
 Dataset Classes Training Validation Testing Task
@@
 ImageNet-R [15] 200 - - 30,000 Robustness of multi-domains
-B.2 Positive and Negative Prompts
-In Table B5, we detail the specific positive and negative prompts utilized for each dataset.
+C.2 Positive and Negative Prompts
+In Table C5, we detail the specific positive and negative prompts utilized for each dataset.
 Additionally, as mentioned in Section 4.1, we incorporate prompts from CuPL [36] to further enhance model performance.
-Table B5: Positive and negative prompts used in experiments. In addition to these prompts, we also employ CuPL [46] prompts to further enhance performance.
+Table C5: Positive and negative prompts used in experiments. In addition to these prompts, we also employ CuPL [36] prompts to further enhance performance.
 Dataset Positive Prompts Negative Prompts
@@
 UCF101 [48] “a photo of a person doing {CLASS}.” “a photo of a person not doing {CLASS}.”
-C Future Work
-In this work, we introduce the first dual-path adapter-style fine-tuning framework for VLMs.
-We believe that the similar concept of dual learning can also be applied to prompt-based learning methods, and be extended to fine-tune other foundational models (e.g., other VLMs [24, 74] and LLMs [63, 5]).
+D License Information
+Datasets. We list the known license information for the datasets below:
+  • MIT License: ImageNet-A [16], ImageNet-V2 [40], ImageNet-R [15], and ImageNet-Sketch [55].
+  • CC BY-SA 4.0 License: OxfordPets [35].
+  • Research purposes only: ImageNet [6], StandfordCars [24], DTD [5], FGVCAircraft [30], SUN397 [63].
+Code. In this work, we also use some code implementations from existing codebase: CLIP [38], CoOp [74], APE [76], and CuPL [36]. The code used in this paper are all under the MIT License.
+E Further Discussions
+Future Work.
+In this work, we introduce negative learning for adapting VLMs to downstream tasks.
+We believe that the similar concept of negative learning can also be applied to prompt-based learning methods, and be extended to fine-tune other foundational models (e.g., other VLMs [21, 64] and LLMs [51, 7]).
 Besides, we also notice that there are a lot of research efforts dedicated to design better prompts (e.g., using LLM [36, 31, 62]) to fully exploit the capabilities of CLIP.
 We hope that with our work, future research endeavors can also be directed to investigate the utilization of negative prompts to better activate the negative inference capabilities, further broadening the scope and effectiveness of CLIP.
+Limitations.
+We identify two potential limitations of our SimNL: (1) Its efficacy in zero-shot situations is constrained due to the scarcity of original negative descriptions in the CLIP training corpus; (2) Similar to other adapter-style fine-tuning approaches, our SimNL fine-tuned on a specific task cannot be directly applied to another task without additional adaptation.
+However, recently, Wang et al.
+[59] shows that adapter-style fine-tuning methods can be extended for these scenarios using the $k$NN algorithm.
+Broader Impacts.
+In this work, we aim to build more reliable machine learning systems by leveraging the extensive knowledge of current foundational models.
+Specifically, we introduce negative learning to more efficiently transfer pre-trained VLMs to downstream tasks, enhancing both the task-specific performance of CLIP and its robustness to natural distribution shifts.
+Additionally, we explore the few-shot adaptation of VLMs in a noisy setting, which better aligns with real-world scenarios where mislabeled samples may exist in the support set.
+We hope this work inspires future studies to focus on the generalization and robustness of pre-trained large-scale foundation models.